### v3

# A large data semi-global existence and convergence theorem for vacuum Einstein’s equations

Abstract.
We prove a semi-global existence and convergence theorem for the $(3+1)$-dimensional vacuum Einstein equations with positive cosmological constant on spacetimes $\widetilde{M}\cong M\times\mathbb{R}$, where $M$ is a closed, connected, oriented three-manifold of negative Yamabe type. In constant mean curvature transported spatial coordinates, we show that solutions arising from a class of arbitrarily large initial data converge to a Riemannian metric of constant negative scalar curvature in infinite Newtonian-like ‘time’. As a consequence, the Einstein-$\Lambda$ flow generically fails to produce geometrization in the sense of Thurston. Our results affirm a conjecture of Ringström concerning the asymptotic in-distinguishability of spatial topology in the large data regime of Einstein-$\Lambda$ dynamics. A related result is established for positive Yamabe type under a technical condition.
1. Introduction
[…]
(1) $\text{Ric}[\widehat{g}]-\frac{1}{2}R[\widehat{g}]\widehat{g}+\Lambda\widehat{g}=0.$
Global hyperbolicity implies the existence of a Cauchy hypersurface and in particular $\widetilde{M}=M\times\mathbb{R}$ with $M$ being diffeomorphic to a Cauchy hypersurface. In this article, we will choose $M$ to be of closed negative Yamabe type (a topological invariant, the so-called $\sigma$ constant $\sigma(M)\leq 0$). By definition, these admit no Riemannian metric $g$ having scalar curvature $R(g)\geq 0$ everywhere. A closed 3-manifold $M$ is of negative Yamabe type if and only if it lies in one of the following three mutually exclusive subsets:
  • (1)
    $M$ is hyperbolizable (that admits a hyperbolic metric),
  • (2)
    $M$ is a non-hyperbolizable $K(\pi,1)$ manifold of non-flat type (the six flat $K(\pi,1)$ manifolds are of zero Yamabe type), here $K(\pi,1)$ is of Eilenberg-MacLane type and by definition $\pi_{1}(M)=\pi$ and $\pi_{i}=0,~i\geq 2$-the universal cover of $K(\pi,1)$ is contractible and known to be diffeomorphic to $\mathbb{R}^{3}$ [44], and
  • (3)
    $M$ has a nontrivial connected sum decomposition (i.e., $M$ is composite)
    (2) $M\approx(\mathsfbf{S}^{3}/\Gamma)_{1}\#..\#(\mathsfbf{S}^{3}/\Gamma)_{k}\#(\mathsfbf{S}^{2}\times\mathsfbf{S}^{1})_{1}\#..\#(\mathsfbf{S}^{2}\times\mathsfbf{S}^{1})_{l}\#\mathsfbf{K}(\pi,1)_{1}\#..\#\mathsfbf{K}(\pi,1)_{m}$
    in which at least one factor is a $K(\pi,1)$ manifold i.e, $m\geq 1$ [46]. In this case the $K(\pi,1)$ factor may be either of flat type or hyperbolizable or non-hyperbolizable non-flat type.
The six flat manifolds comprise by themselves the subset of zero Yamabe type. These admit metrics having vanishing scalar curvature (the flat ones) but no metrics having strictly positive scalar curvature. Here $\Gamma\subset SO(4)$ acts freely and properly discontinuously on $\mathbb{S}^{3}$. Also note that $M\#\mathbb{S}^{3}\approx M$ for any 3-manifold $M$. It is known that every prime $K(\pi,1)$ manifold is decomposable into a (possibly trivial but always finite) collection of (complete, finite volume) hyperbolic and graph manifold  [^2] A compact oriented $3$-manifold $M$ is a graph manifold if $\exists$ a finite disjoint collection of embedded 22-tori $\{T_{j}\}\subset M$ such that each connected component of $M-\cup_{j}T_{j}$ is the total space of circle bundle over surfaces components.
Our goal is to study the long-term behavior of Einstein’s equations 1 on spacetimes $\widetilde{M}=M\times\mathbb{R}$ with $M$ being general 33- manifolds of negative Yamabe type as in 2. An issue closely related to Penrose’s weak cosmic censorship [39] in the cosmological context (compact spatial topology where the notion of null infinity is not defined) may be phrased as follows
Conjecture 1.1 (No-Naked Singularity Conjecture for Cosmological Spacetimes).
[…]
$\operatorname{Ric}_{\widehat{g}}-\tfrac{1}{2}R_{\widehat{g}}\widehat{g}+\Lambda\widehat{g}=\mathcal{T},$
where $\Lambda\in\mathbb{R}$ is the cosmological constant and $\mathcal{T}$ is a smooth energy-momentum tensor obeying the dominant energy condition and representing physically reasonable matter and radiation content.
[…]
Remark 1.
Notice that it is very unlikely to obtain global control of a solution with arbitrary initial data in a fully generic framework since Einstein’s equations are quasi-linear hyperbolic
In this article, we focus on the vacuum case with a positive cosmological constant, i.e., $\mathcal{T}=0$ and $\Lambda>0$.
1.1. Background Literature
The future stability problem for cosmological solutions to the Einstein equations has been extensively studied in the small-data regime, particularly when the spatial manifold MM lies in the negative Yamabe class and admits a hyperbolic metric.
The foundational result in this context is due to Andersson and Moncrief [2], who established the nonlinear future stability of the Milne universe,
$\widehat{g}=-dt^{2}+\frac{t^{2}}{9}\eta,\quad\text{Ric}[\eta]=-\frac{2}{9}\eta,$
on manifolds of the form $\mathbb{H}^{3}/\Gamma\times\mathbb{R}$, where $\Gamma\subset\mathrm{SO}^{+}(1,3)$ is cocompact and acts freely and properly discontinuously on $\mathbb{H}^{3}$. Their result applies to vacuum solutions with spatial topology restricted to eliminate moduli of flat perturbations. The Milne model arises as a quotient of the interior of the future light cone in Minkowski space, and the proof employs a generalized energy method under small perturbations of the hyperbolic metric.
This analysis was extended to higher-dimensional spacetimes ($n+1\geq 4$) in [3], assuming that the spatial manifold admits a strictly stable negative Einstein metric. The proof relies on the decay of an energy functional and the spectral gap of the Lichnerowicz Laplacian associated to the Einstein background.
Subsequent work has incorporated various matter models. Andersson and Fajman [4] established the nonlinear future stability of the Milne model within the Einstein–massive Vlasov system. In a higher-dimensional Kaluza–Klein setting, Branding, Fajman, and Kröncke [24] proved future stability for $(\mathbb{H}^{3}/\Gamma\times\mathbb{T}^{d})\times\mathbb{R}$ with Kaluza–Klein reduction, reducing the system to a coupled Einstein–Maxwell–wave map system.
Further stability results in the small data regime exist for Einstein equations coupled with a positive cosmological constant [13, 12, 34], relativistic fluids [37, 38, 23, 28, 36], and scalar fields [50, 15]. These developments confirm the validity of Conjecture 1.1 for small perturbations of homogeneous background spacetimes with hyperbolic spatial sections.
In contrast, the asymptotically flat case has a distinct lineage, beginning with the monumental work of Christodoulou and Klainerman [10], who proved the global nonlinear stability of Minkowski spacetime. This framework has since been extended in various directions [30, 27, 16, 49, 35], focusing on vacuum and matter models under asymptotically flat conditions.
1.2. Main Result
We consider the Einstein vacuum equations with positive cosmological constant $\Lambda>0$ on $(3+1)$-dimensional globally hyperbolic spacetimes $\widetilde{M}\cong M\times\mathbb{R}$, where $M$ is a closed, connected, oriented $3$-manifold of negative Yamabe type. Importantly, we do not assume that $M$ admits a negative Einstein metric; in particular, $M$ may contain graph manifold summands in its prime decomposition, cf. (2), for which no Einstein metric exists.
In contrast to previous work, we make no smallness assumption on the initial data. We fix the constant mean curvature (CMC) transported spatial gauge. The time parameter is taken to be the mean extrinsic curvature $\tau:=\operatorname{tr}_{g}(k)$, and we consider solutions in the expanding direction, i.e., $\tau<0$ and increasing.
To capture the asymptotic geometry, we introduce a natural geometric rescaling by the function $\varphi^{2}:=\tau^{2}-3\Lambda>0$ (see claim 1) and analyze the rescaled evolution system (64)–(68).
To quantify the deviation of the evolving geometry from hyperbolicity, we define the obstruction tensor
$\mathfrak{T}[g]:=\operatorname{Ric}[g]-\tfrac{1}{3}R(g)g,$
which measures the failure of $g$ to be Einstein. In three spatial dimensions, by Mostow rigidity, $\mathfrak{T}[g]=0$ implies that $g$ is a hyperbolic metric, uniquely determined up to isometry. For generic MM of negative Yamabe type, $\mathfrak{T}[g]\neq 0$ for all $g$.
We impose the transported spatial gauge (see Definition 5), in which the pair $(\Sigma,\mathfrak{T})$ satisfies a manifestly hyperbolic evolution system. This is coupled with an elliptic equation for the lapse function determined by the CMC condition. In this formulation, we establish the following global convergence theorem for the Einstein flow in the large data regime on general negative Yamabe slices.
Theorem 1.1 (Semi-global existence: $\Lambda>0$, $\sigma(M)\leq 0$).
[…]
Fix a smooth background Riemannian metric $\xi_{0}$ on $M$ and a constant $C>1$. For any initial energy quantity $\mathcal{I}^{0}>0$, there exists a constant $a=a(\mathcal{I}^{0})>0$, so that $\mathcal{I}^{0}e^{-a/10}<1$.
Let $(g_{0},\Sigma_{0})$ be an initial data set verifying the Einstein constraint equations at initial CMC time $T_{0}=a>0$, written in CMC-transported spatial coordinates and satisfying:
(3) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},$
(4) $\sum_{I=0}^{3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{I=0}^{2}\left(\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}+\|e^{a}\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\right)\leq\mathcal{I}^{0},$
where $\mathfrak{T}[g_{0}]$ denotes the renormalized trace-free spatial Ricci curvature tensor of $g_{0}$.
Then, the Einstein-$\Lambda$ evolution equations admit a unique classical solution
$T\mapsto(g(T),\Sigma(T))\in\mathcal{C}^{\infty}([T_{0},\infty)\times M)$
in CMC-transported spatial coordinates, satisfying the constraint equations at each slice $T$ and obeying the following uniform a priori estimates for all $T\in[T_{0},\infty)$:
(5) $\sum_{I=0}^{3}\|\nabla^{I}\Sigma(T)\|_{L^{2}(M)}+\sum_{I=0}^{2}\left(\|\nabla^{I}\mathfrak{T}[g(T)]\|_{L^{2}(M)}+\|e^{T}\nabla^{I}\Sigma(T)\|_{L^{2}(M)}\right)\leq C_{1}(1+\mathcal{I}^{0}),$
(6) $C_{2}^{-1}g_{0}\leq g(T)\leq C_{2}g_{0}.$
Here, $C_{1},C_{2}>0$ are numerical constants depending only on the universal geometric and analytic data of the problem (e.g., Sobolev constants of $(M,g_{0})$ and the constants in the structure equations, dimension of $M$), but independent of $T$. The developed spacetime is future geodesically complete.
Moreover, the solution $(g(T),\Sigma(T))$ converges in the $C^{\infty}$ topology, as $T\to\infty$, to a limiting Riemannian metric $\widetilde{g}$ on MM of pointwise constant negative scalar curvature, in the sense that
$\Sigma(T)\to 0\quad\text{and}\quad g(T)\to\widetilde{g},\quad\text{as }T\to\infty,$
[…]
Remark 2.
Note that the allowed initial data size $\mathcal{I}^{0}$ is modulated by the condition $\mathcal{I}^{0}e^{-a/10}<1$ for initial time $T_{0}=a>0$ i.e., when $a$ is small, then the allowed data size is $O(1)$ but when $a\gg 1$, $\mathcal{I}^{0}$ can be very large.
Remark 3.
The result of the convergence of an arbitrary metric to the constant negative scalar curvature metric can be compared to the Yamabe flow on negative Yamabe manifolds. Of course, the convergence of the Yamabe flow on manifolds with $\sigma(M)\leq 0$ is trivial compared to its positive counterpart. Contrary to the Einstein-$\Lambda$ flow, the Yamabe flow is of energy-critical parabolic nature and endowed with monotonic entities (e.g., similarly, Yang-Mills heat flow [51] is of parabolic nature). Moreover, the limit constant scalar curvature metric obtained in the framework of Einstein’s equations is not conformal to the initial metric.
Remark 4.
This reproduces the elementary fact that in dimension $3$, there is an infinite-dimensional family of constant negative scalar curvature metrics that are non-conformal and non-isometric to the standard metrics on closed hyperbolic manifolds. Note that standard metrics are scalar curvature rigid in the conformal class i.e., conformal deformations of finite volume hyperbolic metrics preserving scalar curvature are necessarily isometries.
1.3. Heuristics on the genericity of the initial data
We comment on the notion of “genericity” of the initial data from a heuristic yet technically relevant viewpoint. The compact spatial manifold MM under consideration is of negative Yamabe type and, in particular, admits no Riemannian metric with nontrivial global Killing fields. Indeed, any compact 33–manifold that supports a continuous isometry group necessarily admits an effective SO(2)SO(2)–action. The classification theorem of Fischer–Moncrief [18] implies that no nontrivial connected sum of negative Yamabe type supports such an action. Among prime manifolds, only a distinguished subclass of Seifert fibered spaces—specifically those whose fundamental group has nontrivial infinite cyclic center—can admit an $SO(2)$ symmetry. Even in these exceptional geometries, such continuous symmetries arise only in a nongeneric subset of the moduli space of initial data. Thus, from the perspective of geometric constraints, the absence of continuous symmetries is the generic situation.
Although the topology of $M$ allows for considerable freedom in the selection of initial data, the data are not arbitrary from the viewpoint of their quantitative concentration properties. To analyze this, we introduce a heuristic amplitude–wavelength characterization of the pair $(\Sigma,\mathfrak{T})$, where $\Sigma$ denotes the trace–free part of the second fundamental form and $\mathfrak{T}$ denotes the obstruction tensor. We let
$(A_{\Sigma},L_{\Sigma}),\qquad(A_{\mathfrak{T}},L_{\mathfrak{T}})$
denote the respective amplitude–wavelength pairs of $\Sigma$ and $\mathfrak{T}$. Interpreting $(\Sigma,\mathfrak{T})$ as wave-type fields, we view $L_{\Sigma}$ and $L_{\mathfrak{T}}$ as characteristic length scales, while $A_{\Sigma}$ and $A_{\mathfrak{T}}$ denote dimensionally rescaled amplitudes.
The control of the initial data norm in Theorem 1.1 implies, at the highest order, the constraint
(7) $L_{\Sigma}^{-3}|A_{\Sigma}|^{2}\;+\;L_{\Sigma}^{-1}|A_{\Sigma}|^{2}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}^{3}\;\lesssim\;\big(\mathcal{I}^{0}\big)^{2},$
whereas for the lower-order norms one obtains
(8) $L_{\Sigma}^{-1}|A_{\Sigma}|^{2}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}^{3}\;\lesssim\;e^{-2a}\big(\mathcal{I}^{0}\big)^{2}.$
Analogously, for the obstruction tensor $\mathfrak{T}$,
(9) $L_{\mathfrak{T}}^{-1}|A_{\mathfrak{T}}|^{2}\;+\;|A_{\mathfrak{T}}|^{2}L_{\mathfrak{T}}\;+\;|A_{\mathfrak{T}}|^{2}L_{\mathfrak{T}}^{3}\;\lesssim\;\big(\mathcal{I}^{0}\big)^{2},$
and by the relative smallness condition stated in Theorem 1.1, one may in particular have
$\mathcal{I}^{0}=O\!\left(e^{\frac{a}{10}}\right).$
Observe that the presence of the term $L_{\Sigma}^{-3}|A_{\Sigma}|^{2}$ in (7) allows for substantial concentration provided the lower-order bound (8) remains satisfied. For instance, if
$A_{\Sigma}=O\!\left(e^{-\alpha a}\right),\qquad L_{\Sigma}=O\!\left(e^{-\beta a}\right),\qquad\text{with}\quad 2\alpha+3\beta=\tfrac{9}{5},$
then for $a>0$, the inequalities (7)–(8) are compatible. Thus, for $\Sigma$, high-frequency small-amplitude concentration is allowed, whereas large amplitude on very small scales is prohibited. In sharp contrast, (9) demonstrates that $\mathfrak{T}$ is significantly more flexible: one may heuristically take
$|A_{\mathfrak{T}}|^{2}=O\!\left(e^{\tfrac{a}{10}}\right),\qquad L_{\mathfrak{T}}=O\!\left(e^{-\tfrac{a}{10}}\right),$
so that large amplitude concentrated on small scales is admissible for $\mathfrak{T}$ when $a\gg 1$.
To interpret this from the viewpoint of free gravitational radiation, let us decompose the Weyl curvature into its electric and magnetic components relative to the unit timelike vector $T$:
$E:=W(T,\cdot,T,\cdot),\qquad B:={}^{*}W(T,\cdot,T,\cdot).$
From the Gauss equation,
$E=\mathfrak{T}+O\big(|\Sigma|^{2}\big),$
and hence one may allow initial data with
$|A_{E}|^{2}=O\!\left(e^{\tfrac{a}{10}}\right),\qquad L_{E}=O\!\left(e^{-\tfrac{a}{10}}\right),$
which corresponds to high-frequency, large-amplitude electric Weyl curvature (short-pulse type) even in the positive cosmological constant regime. By contrast, via the Codazzi identity,
$B=O\big(\nabla\Sigma\big),$
the magnetic part inherits the amplitude–wavelength restrictions of $\Sigma$, and thus only high-frequency, small-amplitude configurations are allowed for $B$ at the initial slice. See Figure 1 for a schematic illustration.
[…]
$|A_{B}|=O\!\left(e^{\tfrac{a}{20}}\right),\qquad L_{B}=O\!\left(e^{-\tfrac{a}{10}}\right),$
then gravitational collapse could ensue despite the rapid expansion induced by the positive cosmological constant. This suggests that the class of initial data considered here carries genuine physical significance, rather than merely constituting a nonempty technical scenario. A detailed semi-global existence analysis, including the potential formation of trapped surfaces in the cosmological regime, is currently in progress.
Figure 1. A rough heuristic depicting the characterization of the smooth initial data considered in the article. The horizontal axis denotes the support size on the initial slice $(M,g_{0})$ while the vertical axis does so for the amplitude. $a$ is the free parameter considered here that can modulate the necessary amplitude-wavelength as to how big the initial data can be, dependent on the $CMC$ time $T_{0}=a$ of the initial Cauchy slice. While the electric part of spacetime Weyl curvature can have any data of the three types including the data of short pulse type (type 2) shown above, the magnetic part of the same cannot be of the short pulse type. Type (3) is a typical magnetic field (of spacetime Weyl curvature) configuration that is allowed in this study. Here, the initial re-scaled geometry $(M,g_{0})$ is assumed to have $O(1)$ diameter.
A corollary to the theorem 1.1 is as follows
Corollary 1.1.
Conjecture 1.1 holds in the class of $(3+1)$-dimensional solutions to the vacuum Einstein equations with positive cosmological constant $\Lambda>0$, defined on globally hyperbolic spacetimes $\widetilde{M}\cong M\times\mathbb{R}$ where $M$ is a closed $3$-manifold of negative Yamabe type. The result holds for a class of large initial data prescribed on a Cauchy hypersurface at a Newtonian-like time $T_{0}=a$. In particular, no future naked singularities can form within this setting.
1.4. Comparison with previous studies
The seminal heuristic investigation of the asymptotic dynamics of initially expanding spatially homogeneous cosmological spacetimes governed by Einstein’s field equations with a positive cosmological constant $\Lambda>0$ was undertaken by Wald [48], wherein it was suggested that such solutions generically exhibit de Sitter-like behavior at late times. Subsequently, Ringström [29] established a rigorous framework for analyzing the future development of spatially homogeneous solutions by localizing the global Cauchy problem to a suitable coordinate neighborhood. Exploiting this localization, he proved nonlinear stability and future geodesic completeness for broad classes of spatially homogeneous cosmological models with arbitrary compact spatial topologies.
[…]
(10) $\widetilde{g}=\Omega^{2}g,$
supplemented by a hierarchy of conformal curvature variables and auxiliary fields required to ensure hyperbolicity and regularity at the conformal boundary. The initial data are posed not on a standard Cauchy hypersurface but rather at past or future conformal infinity $\mathcal{J}^{\pm}$, where one prescribes a Riemannian metric $\widehat{h}_{ij}$ on the 33–manifold $\mathcal{J}^{\pm}$ (defined only up to an overall conformal scaling) together with a trace–free symmetric 22–tensor $\widehat{d}_{ij}$, subject solely to a divergence constraint with respect to $\widehat{h}$. Crucially, in contrast to the Cauchy formulation, no Hamiltonian (scalar) constraint is imposed at $\mathcal{J}^{\pm}$; the conformal transformation effectively removes the obstruction associated with the scalar constraint at the hypersurface on which $\Omega=0$.
[…]
By contrast, the present work addresses a fundamentally distinct regime: it employs a CMC–time slicing on a spatially compact spacetime. The construction accommodates genuinely large data and is intrinsically tied to geometric control of high-order norms within the CMC foliation, rather than to asymptotic smallness near $\mathcal{J}^{\pm}$. In particular, the Hamiltonian and momentum constraints are fully retained and enforced at the initial Cauchy hypersurface, in stark contrast with Friedrich’s boundary data prescription. Furthermore, in the absence of $\Lambda>0$, the conformal extension techniques utilized by Friedrich are, in general, obstructed by the nonexistence of smooth conformal completions of asymptotically simple solutions to the pure vacuum system $\mathrm{Ric}(g)=0$. This obstruction highlights the structural incompatibility between conformal compactification and the non–conformally rescaled large–data interior approach developed in the present analysis.
In contrast, the present study is formulated on a physical compact Cauchy slice at large CMC–time $T_{0}=a$, of negative Yamabe type. We work with the full set of Cauchy data $(g_{0},k_{0})$ satisfying the nonlinear Einstein constraint equations. Furthermore, to the author’s knowledge, Friedrich’s results pertain to an open set of data near de Sitter, typically on manifolds of spherical topology, whereas our construction admits large data consistent with the prescribed CMC hierarchy. In particular, the initial profiles may exhibit strongly concentrated gravitational radiation (reasonably high-frequency), with amplitudes of order $O(e^{a/10})$ supported in regions of spatial extent $O(e^{-a/10})$, while still obeying the required geometric and analytic constraints.
[…]
In the present work, we extend this line of inquiry in several significant directions. First, we dispense with the assumption of a homogeneous background solution and instead consider a large class of initial data sets satisfying the Einstein constraint equations on compact manifolds MM of negative Yamabe type, specifically those with $\sigma(M)\leq 0$, as characterized in equation (2). These manifolds rarely admit any locally homogeneous (or isotropic) metrics. Our main result constitutes, to the best of our knowledge, the first rigorous demonstration of the global-in-time well-posedness of solutions to the Einstein equations with $\Lambda>0$, emanating from large, non-perturbative, and reasonably generic initial data on compact manifolds with $\sigma(M)\leq 0$ in dimension $n=3$.
1.5. Novelty and the main difficulty
In the analysis of nonlinear hyperbolic evolution equations, it is often heuristically observed that small data global existence on $0,\infty)$ may correspond to weighted large data global existence on $T_{0},\infty)$ for sufficiently large $T_{0}$. In the context of Einstein’s equations with positive cosmological constant, one might conjecture that the small data global existence results established in [13, 34] could be extended to yield large data (unweighted) results by considering initial time $T_{0}\gg 1$. This expectation is, however, not valid in the current setting for both geometric and analytic reasons.
To elucidate this failure of transference, consider initial data at time $T_{0}\gg 1$ satisfying
(11) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},\qquad\sum_{0\leq I\leq 2}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\leq\epsilon_{0},$
along with high-order control:
(12) $\sum_{0\leq I\leq 3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{0\leq I\leq 2}\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}\leq\mathcal{I}^{0},$
where $\mathfrak{T}[g]:=\operatorname{Ric}[g]-\frac{1}{n}R(g)g$ is the trace-free Ricci tensor. $\epsilon_{0}$ is sufficiently small and $\mathcal{I}^{0}$ is sufficiently large. Despite the smallness of $\Sigma_{0}$ in low-order norms, the high-order norms of $\Sigma_{0}$ and all norms of $\mathfrak{T}[g_{0}]$ may remain large. As a consequence, the data do not satisfy the hypotheses of small data stability results such as those in [13, 34], and no global control follows.
By contrast, suppose that at $T_{0}=0$, both $\Sigma_{0}$ and $\mathfrak{T}[g_{0}]$ satisfy uniform smallness conditions:
(13) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},\qquad\sum_{0\leq I\leq 2}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\leq\epsilon_{0},$
(14) $\sum_{0\leq I\leq 3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{0\leq I\leq 2}\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}\leq\epsilon_{0},$
for $\epsilon_{0}$ sufficiently small. Then the global existence and nonlinear stability results of [13] are applicable. However, it is a consequence of [52] that these assumptions (along with a lower bound on the eigenvalue of the Lichnerowicz Laplacian) imply that the initial Cauchy hypersurface $(M,g_{0})$ admits an Einstein metric, i.e., $(M,g_{0})$ is hyperbolizable when $n=3$. Thus, the result becomes restricted to small perturbations of spatially homogeneous (e.g., de Sitter) solutions and does not extend to general negative Yamabe manifolds, such as those with graph manifold structure. Therefore, this case can never yield any global control in the previous case when the obstruction tensor $\mathfrak{T}$ and the top order norm of $\Sigma_{0}$ are large. In more physical terms, as stated in section 1.2, the data is large in the sense that the electric part of the Weyl curvature (free gravity) can be arbitrarily large and concentrated on small measure subsets. This is novel and to our knowledge the first result of this kind.
In contrast, the present article develops a framework in which the initial spatial metric $g_{0}$ is arbitrary (subject to geometric bounds), and $\mathfrak{T}[g_{0}]$ is bounded but not required to be small. No Einstein metric structure on the Cauchy hypersurface is assumed. This generality permits global existence and control of large data on non-hyperbolizable topologies, significantly extending the scope of prior stability results.
Secondly, observe that the Hamiltonian constraint associated with the Einstein equations implies that the scalar curvature of the spatial metric $g$ satisfies
(15) $R(g)+\frac{n-1}{n}=|\Sigma|^{2}.$
Under the hypotheses of Theorem 1.1, in particular the bound on $\Sigma$ in $W^{2,2}$, it follows immediately from (15) that
(16) $\left\|R(g)+\frac{n-1}{n}\right\|_{W^{2,2}(M)}\leq e^{-a}\mathcal{I}^{0}.$
For $a\gg 1$ and $\mathcal{I}^{0}$ fixed, or $\mathcal{I}^{0}$ large but $a$ sufficiently large in terms of $\mathcal{I}^{0}$, the right-hand side can be made small. However, this does not imply that $g$ is close to a metric of constant scalar curvature in a strong topology.
Indeed, for $n\geq 3$, the mapping $g\mapsto R(g)$ is a nonlinear second-order differential operator. More precisely, closeness of scalar curvature in $W^{2,2}$ to a negative constant does not control the closeness of the metric $g$ to a reference constant scalar curvature metric $\gamma$ in, say, $W^{4,2}$ or stronger norms. Consequently, even though $R(g)\approx-\frac{n-1}{n}$ in $W^{2,2}$, the metric $g$ itself may be arbitrarily far from any metric of constant scalar curvature $\frac{n-1}{n}$ in any Sobolev norms.
Furthermore, the bounds in Theorem 1.1 only require $g$ to satisfy uniform ellipticity and boundedness of the trace-modified Ricci tensor $\mathfrak{T}[g]:=\operatorname{Ric}[g]-\frac{1}{n}R(g)g$ in $W^{2,2}(M)$, and thus permit large deviations from any background Einstein metric. Therefore, the convergence result established in Theorem 1.1 describes a genuinely large data dynamical regime, in which the initial metric $g_{0}$ may be far in $W^{4,2}(M)$ (and higher topologies) from any constant scalar curvature or Einstein geometry (even if it exists on $M$). Nevertheless, the solution converges globally in time to a limiting metric of constant scalar curvature.
This phenomenon—global convergence to a constant scalar curvature geometry from arbitrary (non-small) initial data—is not captured by previous small perturbation results around Einstein metrics (cf. [13]). To the best of the author’s knowledge, this work constitutes the first rigorous treatment of the large data convergence problem for the Einstein equations with positive cosmological constant on general compact spatial manifolds in dimensions $n\geq 3$.
We now elaborate on a key structural aspect of our formulation. Rather than employing the coupled hyperbolic-elliptic system for $(g,\Sigma)$ in constant mean curvature (CMC) and spatial harmonic (SH) gauge, we adopt the manifestly hyperbolic evolution system for $(\Sigma,\mathfrak{T})$, where $\mathfrak{T}_{ij}:=\operatorname{Ric}_{ij}-\frac{1}{n}R(g)g_{ij}$, coupled to an elliptic lapse equation (76), and impose the CMC-transported spatial coordinate gauge (i.e., vanishing shift vector field). This formulation is motivated by the classical work of Choquet-Bruhat and York [1], who studied a hyperbolic system for $(\Sigma,\operatorname{Ric})$. In this gauge, the evolution equation for $g$ (cf. (64)) is treated as a transport equation along the future-directed unit normal vector field $\partial_{T}$, whose flow lines form the characteristics. In contrast, under spatial harmonic gauge, the metric evolution is governed by a quasilinear wave equation with null cone characteristics. More significantly, the SH gauge introduces serious analytical complications in the large data regime, which we now make precise.
Preservation in time of the SH gauge condition $V^{i}=0$, where $V^{i}:=g^{jk}(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk})$ is the tension field associated to a fixed background metric $\xi$, yields the following elliptic equation for the shift vector field $X$:
(17) $\mathcal{P}X^{i}:=-\Delta_{g}X^{i}-\operatorname{Ric}^{i}_{j}[g]X^{j}+2\nabla^{j}X^{k}(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk})$ $=\mathcal{F}^{i}[g,\Sigma,N],$
where $\mathcal{F}^{i}[g,\Sigma,N]$ denotes a source term depending on the lapse $N$, the second fundamental form $\Sigma$, and $g$ relative to $\xi$ (cf. [13, 34]). In previous small data studies [13, 34], $\xi$ is chosen to be a fixed hyperbolic metric and $g$ is assumed close to $\xi$ in $W^{s,2}$ with $s>\frac{5}{2}$, so that $\mathcal{P}$ defines an isomorphism $W^{s+1,2}\to W^{s-1,2}$.
Let us sketch the standard argument in the small data regime. The operator $\mathcal{P}$ is elliptic and may be decomposed as
(18) $\mathcal{P}X^{i}+\mathscr{L}_{X}V^{i}$ $=\Delta^{\xi}_{g}X^{i}-g^{k\ell}R[\xi]^{i}{}_{kj\ell}X^{j},$
where $\mathscr{L}_{X}V$ denotes the Lie derivative of the tension field, and $\Delta^{\xi}_{g}$ is the rough Laplacian defined using $\xi$ and acting on vector fields (cf.  [3]). The derivation follows from differentiating the pullback identity
$(\psi_{s}^{*}V)^{i}=(\psi_{s}^{*}g^{-1})^{jk}(\Gamma[\psi_{s}^{*}g]^{i}_{jk}-\Gamma[\psi_{s}^{*}\xi]^{i}_{jk}),$
where $\psi_{s}$ is the flow of $X$. In particular, under the smallness assumption $\|g-\xi\|_{W^{s,2}}\ll 1$, the operator $\mathcal{P}_{g,\xi}:=\Delta^{\xi}_{g}-g^{k\ell}R[\xi]^{i}{}_{kj\ell}$ becomes coercive. Indeed, integration by parts yields:
$0=\int_{M}\xi_{ij}Z^{i}\mathcal{P}_{g,\xi}Z^{j}\,d\mu_{\xi}$ $=\int_{M}\left[\xi_{ij}\xi^{k\ell}\nabla_{k}Z^{j}\nabla_{\ell}Z^{i}-\xi_{im}\xi^{k\ell}Z^{m}R[\xi]^{i}{}_{kj\ell}Z^{j}+\text{Err}_{g,\xi}\right]d\mu_{\xi},$
where the error terms $\text{Err}_{g,\xi}$ are perturbative if $\|g-\xi\|_{W^{s,2}}$ is small. Since $R[\xi]^{i}{}_{j}$ is negative-definite (in the case $\xi$ is hyperbolic), the integrand is strictly positive unless $Z\equiv 0$, hence $\ker\mathcal{P}_{g,\xi}=\{0\}$ and the operator is an isomorphism (note it is self-adjoint and elliptic).
However, this strategy fails in the large data regime. In particular, the general Cauchy hypersurfaces considered here are compact, but may carry no hyperbolic metric at all, for example when $\sigma(M)\leq 0$. Moreover, in the absence of smallness assumptions on $\|g-\xi\|_{W^{s,2}}$, the leading-order perturbative structure in the elliptic operator $\mathcal{P}$ degenerates, and both the Ricci and first-order connection terms in (17) are no longer controlled. This introduces Gribov-type ambiguities in solving for $X$, akin to gauge degeneracy in non-Abelian Yang-Mills theory in Coulomb gauge [18]. Indeed, the SH gauge is analogous to the Coulomb gauge in Yang–Mills theory, where existence and uniqueness of global solutions is obstructed by the presence of Gribov horizons.
In contrast, the CMC-transported spatial coordinate gauge circumvents this obstruction entirely: the shift vector field is identically zero, and the evolution system is strictly hyperbolic modulo a single elliptic lapse equation. This choice parallels the temporal gauge in Yang–Mills theory, which avoids the Gribov ambiguity but introduces technical subtleties in constraint propagation. Our analysis reveals that in the gravitational setting, the CMC-transported spatial gauge provides a natural framework for addressing large data evolution problems without encountering elliptic degeneracy in gauge fixing.
A key technical ingredient in the analysis is the construction of a hierarchy of energy norms for the transverse-traceless second fundamental form $\Sigma$. These norms are designed to exploit the decay induced by the cosmological expansion. Specifically, although the top-order energy norm of $\Sigma$ does not decay (as reflected in Theorem 1.1), the lower-order norms exhibit exponential decay, which is systematically leveraged in a bootstrap framework to gain successive improvements at each level of regularity. This hierarchy is essential for closing the top-order estimates.
The structure of the Einstein equations plays a crucial role in controlling nonlinear interactions. In particular, the terms that lack decay at top order always appear with an exponentially decaying prefactor $e^{-T}$. While this is not a manifestation of a null structure, it indicates a hidden weak null-type behavior that remains to be fully understood. Furthermore, although the normalized lapse perturbation $\frac{N}{n}-1$ does not decay at top order, all instances where this quantity appears differentiated at the top order are accompanied by a decaying factor $e^{-T}$, while in undamped terms it arises only through lower order derivatives, which ensures integrable decay. We elaborate on this more in section 1.6
All energy and elliptic norms are defined relative to the evolving spatial metric $g$, and hence a priori estimates for $g$ are indispensable. To control the metric, we integrate the evolution equation
(19) $\partial_{T}g_{ij}=-\frac{2\varphi}{\tau}N\Sigma_{ij}-2\left(1-\frac{N}{n}\right)g_{ij},$
along the CMC-transported foliation. As $\Sigma$ is controlled in $W^{I,2}$, this yields control over $g$ in $W^{I,2}$, but not in $W^{I+1,2}$. Consequently, standard elliptic regularity theory, which would require $g\in W^{I+1,2}$ to control $\frac{N}{n}-1$ in $W^{I+2,2}$, does not directly apply.
This leads to an apparent regularity obstruction in estimating the normalized lapse from its governing elliptic equation:
(20) $\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|^{2}+\frac{1}{n}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|^{2}.$
If Sobolev norms were defined relative to a fixed background metric $\xi_{0}$, as in [3], the operator on the left-hand side would define an isomorphism between appropriate Sobolev spaces, yielding the estimate
(21) $\left\|\frac{N}{n}-1\right\|_{W^{I+2,2}(M)}\lesssim\|\Sigma\|^{2}_{W^{I,2}(M)},\qquad I>\tfrac{3}{2}.$
However, in the present setting, we do not have access to the necessary $W^{I+1,2}$ regularity of $g$, either via transport or elliptic estimates, due to the absence of spatial harmonic gauge.
To circumvent this, we define Sobolev norms with respect to the dynamical metric $g$ and commute the elliptic equation (20) with covariant derivatives associated to the Levi-Civita connection of $g$. Upon commuting, curvature terms arise via standard identities, and elliptic regularity for (20) then requires control over the $W^{I-1,2}$ norm of the Riemann curvature of $g$ to estimate $\frac{N}{n}-1$ in $W^{I+2,2}$. One subtlety here is that we need to control the metric components separately through transport equation.
This is resolved through the algebraic decomposition of the Riemann tensor in terms of the electric Weyl tensor $\mathfrak{T}$ and scalar curvature $R[g]$. The latter is in turn expressed via the Hamiltonian constraint:
(22) $R[g]+\frac{n-1}{n}=|\Sigma|^{2}.$
Hence, the Riemann tensor is controlled in $W^{I-1,2}$ in terms of the $W^{I-1,2}$ norms of $\Sigma$ and $\mathfrak{T}$. This suffices to control all curvature-dependent commutator terms in the elliptic theory.
Finally, since the wave equations require $W^{I,2}$ control of $\Sigma$ and $W^{I-1,2}$ control of $\mathfrak{T}$, the above argument provides the needed $W^{I-1,2}$ regularity of the Riemann tensor. Consequently, elliptic regularity theory yields (21), thereby closing the necessary estimates for the lapse.
A further distinguishing feature of this work is the generality of the class of spatial topologies to which our analysis applies. Specifically, we establish our semi-global well-posedness result for every smooth, closed, connected, oriented 3-manifold that admits a metric of negative Yamabe type, i.e., one for which no conformal class contains a metric of non-negative scalar curvature. Equivalently, this includes all manifolds falling into Classes 1–3 in the standard topological classification of closed 3-manifolds. To our knowledge, no prior analysis of the full Einstein equations—particularly at the level of large data energy estimates—has been carried out for such a topologically general class of Cauchy hypersurfaces. Lastly, a major novel feature of this study is sketched out in section 1.3.
[…]
1.6. Main Idea of the Proof
From a structural standpoint, the Einstein evolution in vacuum (with vanishing cosmological constant $\Lambda=0$), when expressed in a gauge adapted to the underlying geometric dynamics, constitutes a quasi-linear hyperbolic system subject to elliptic constraints. In particular, unlike the Ricci flow [40, 41, 42], which is manifestly parabolic, or even the semi-linear Yang–Mills heat flow [51], the Einstein equations do not admit any global-in-time regular enough coercive functional of geometric origin that could, in principle, act as a Lyapunov quantity. The absence of such a monotone or dissipative functional precludes, in general, any a priori global control of the evolving geometry beyond the small-data regime.
[…]
(23) $\frac{\partial g_{ij}}{\partial T}$ $=-2N\Sigma_{ij}-2\Bigl(1-\frac{N}{n}\Bigr)g_{ij}-(\mathscr{L}_{X}g)_{ij},$
$\frac{\partial\Sigma_{ij}}{\partial T}$ $=-(n-1)\Sigma_{ij}-N\Bigl(\mathrm{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij}\Bigr)+\nabla_{i}\nabla_{j}\Bigl(\frac{N}{n}-1\Bigr)$
(24) $\quad+2N\Sigma_{ik}\Sigma^{k}_{\;j}-\frac{1}{n}\Bigl(\frac{N}{n}-1\Bigr)g_{ij}-(n-2)\Bigl(\frac{N}{n}-1\Bigr)\Sigma_{ij}-(\mathscr{L}_{X}\Sigma)_{ij},$
(25) $R+\frac{n-1}{n}-|\Sigma|^{2}$ $=0,$
(26) $\nabla_{j}\Sigma^{j}_{\;i}$ $=0.$
In this gauge the spatial operator
$g\longmapsto\mathrm{Ric}[g]+\frac{n-1}{n^{2}}g$
is elliptic. Thus, the system (23)–(68) is hyperbolic in its principal part, augmented by elliptic equations controlling the gauge variables $(N,X)$. The evolution of $\Sigma$ is governed by two competing mechanisms:
  • a linear dissipative term $(n-1)\Sigma_{ij}$, reflecting the intrinsic expansion of slices with negative scalar curvature;
  • a quadratic Riccati-type interaction $2N\Sigma_{ik}\Sigma^{k}_{\;j}$, which tends to induce growth.
Additional nonlinear contributions arise through the curvature term $\mathrm{Ric}_{ij}$ and, more subtly, via the lapse deformation factor (N/n)1(N/n)-1, whose control is itself quadratic in $\Sigma$.
To seek control, one may impose a bootstrap assumption asserting smallness of $\Sigma$ and $\mathfrak{T}:=\mathrm{Ric}[g]-\frac{1}{n}\text{Scal}[g]g$ in appropriate Sobolev (scale-invariant) norms and then derive energy estimates for a suitably corrected energy functional (as constructed in [3]) applied to the pair $(\Sigma,\mathfrak{T})$. An alternative construction considers weighted quantities of the form $(e^{\alpha T}\Sigma,e^{\alpha T}\mathfrak{T})$, for a parameter $\alpha\in(0,n-1)$, coupled with initial data prescribed at a large time $T_{0}=a$. Observe, however, that
$e^{\alpha T}\|(\Sigma,\mathfrak{T})\|\leq\mathcal{I}\quad\text{at}\quad T=T_{0}=a\gg 1$
is equivalent to the smallness, in a scale-free norm, of the unweighted fields $(\Sigma,\mathfrak{T})$. Thus, both approaches reduce essentially to small-data analysis.
Consequently, in the case $\Lambda=0$, no global control for large initial data is attainable via currently available techniques, and in fact such control is not expected to hold in general. In the small-data regime of [3], the linear damping dominates the nonlinear terms, yielding global existence and asymptotic stability. In contrast, when the initial data are large, the quadratic Riccati-type growth overtakes the linear decay, rendering the higher-order geometric quantities uncontrollable over long times. This intrinsic conflict between dissipative and self-amplifying contributions—exacerbated by the quadratic dependence of the lapse deformation on the fields—explains the fundamental obstruction to large-data global-in-time control in the hyperbolic Einstein evolution.
When $\Lambda>0$, the structure of the Einstein evolution equations acquires a qualitatively and quantitatively improved feature. Specifically, all nonlinear Riccati-type contributions and curvature couplings of the form
$\mathfrak{T}_{ij}:=\mathrm{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij}$
appear multiplied by a universal exponentially decaying factor $e^{-T}$. Schematically, the evolution of $\Sigma_{ij}$ may be expressed as
$\frac{\partial\Sigma_{ij}}{\partial T}$ $\approx-(n-1)\Sigma_{ij}-e^{-T}N\Bigl(\mathrm{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij}\Bigr)+e^{-T}\nabla_{i}\nabla_{j}\Bigl(\frac{N}{n}-1\Bigr)+2e^{-T}N\Sigma_{ik}\Sigma^{k}_{\;j}$
(27) $\quad-\frac{1}{n}e^{-T}\Bigl(\frac{N}{n}-1\Bigr)g_{ij}-(n-2)e^{-T}\Bigl(\frac{N}{n}-1\Bigr)\Sigma_{ij}.$
The exponential damping factor $e^{-T}$ acts as a geometric suppression of nonlinear interactions, which we entitle as the weak null structure in Einstein-$\Lambda$ setting. Recall that for Minkowski space $(\mathbb{R}^{3+1},\eta)$, the null condition of Klainerman [25] imposes algebraic cancellation in quadratic nonlinear terms of the schematic form
$\Box_{\eta}\varphi=\mathcal{N}(\partial\varphi,\partial\varphi),$
ensuring that no product of two “bad” derivatives occurs in characteristic directions. This mechanism prevents the focusing of nonlinear effects along null hypersurfaces and underlies small-data global existence results. More refined versions of such structures, termed weak null conditions, arise for quasi-linear wave equations, notably in the geometric spacetime harmonic gauge formulation of the Einstein vacuum equations [30].
In the present setting, the term $2N\Sigma_{ik}\Sigma^{k}_{\;j}$ plays an analogous role to the nonlinearity $(\partial\varphi)^{2}$ in the wave equation. Without suppression, this Riccati-type interaction would typically lead to finite-time blow-up. However, due to the presence of the exponentially decaying prefactor $e^{-T}$, arising from the positive cosmological constant, the effective strength of the nonlinearity decreases along the evolution, thereby enabling controlled long-time propagation, even for initial data of large amplitude, provided that the initial time slice is chosen at sufficiently large $T_{0}$.
Furthermore, the curvature obstruction tensor $\mathfrak{T}$ also appears multiplied by $e^{-T}$, which permits direct control of the lower-order norms of $\Sigma$ via integration of equation (27). Notably, such control remains valid without requiring smallness assumptions on $\mathfrak{T}$. This should be contrasted with the vacuum case $\Lambda=0$, cf. (24), where no such damping is present and hence uniform-in-time control of $\Sigma$ cannot be established without imposing smallness on $\mathfrak{T}$ at the outset. Let us illustrate the heuristics.
For $\Lambda>0$, the obstruction tensor $\mathfrak{T}$ approximately satisfies
(28) $\frac{\partial\mathfrak{T}}{\partial T}\approx Ne^{-T}\big(\Delta_{g}\Sigma+\mathrm{Riem}\cdot\Sigma\big)+e^{-T}\nabla^{2}\!\left(\frac{N}{n}-1\right)+e^{-T}\nabla\!\left(\frac{N}{n}-1\right)\Sigma$
(29) $+e^{-T}\mathfrak{T}\cdot\Sigma+\left(\frac{N}{n}-1\right)g.$
In the constant-mean-curvature (CMC) transported spatial gauge ($X=0$), the system of equations (27-28) form a hyperbolic system, supplemented by an elliptic equation for the lapse function $N$. For the sake of exposition, we assume that $N$ can be solved smoothly from this elliptic equation (though the analytic subtleties involved are addressed elsewhere in the introduction).
The analytical difficulty in the $\Lambda>0$ case is to achieve uniform global control of the geometry for all $T\in[T_{0},T_{\infty}]$. In the rescaled setting, this amounts to proving that curvature, deformation, and gauge variables remain uniformly bounded for all future times. To measure these quantities precisely, we introduce three scale-invariant norms:
$\mathcal{O}:=\sum_{I\leq 3}\Big[\|\nabla^{I}\Sigma\|_{L^{2}(M)}+\|\nabla^{I-1}\mathfrak{T}\|_{L^{2}(M)}\Big],\quad\mathcal{F}:=\sum_{I\leq 2}\|e^{\gamma T}\nabla^{I}\Sigma\|_{L^{2}(M)},$
$\mathcal{N}^{\infty}:=\|e^{\gamma T}\Sigma\|_{L^{\infty}(M)}+\Big\|e^{2\gamma T}\!\Big(\frac{N}{n}-1\Big)\Big\|_{L^{\infty}(M)},\qquad\gamma\in(0,1].$
Heuristically:
  • $\mathcal{O}$ controls the curvature energy of the spacetime;
  • $\mathcal{F}$ measures a renormalized “entropy” (reverse) associated to the transverse-traceless part of $\Sigma$;
  • $\mathcal{N}^{\infty}$ quantifies the pointwise amplitude of the gauge and deformation variables.
Together, they form a closed hierarchy of control for the full system.
The proof proceeds by a bootstrap argument, standard in nonlinear geometric analysis. We assume that on some interval $T_{0},T]\subset[T_{0},T_{\infty})$, the following bounds hold:
$\mathcal{O}(T)\leq\Gamma,\qquad\mathcal{F}(T)\leq\mathds{Y},\qquad\mathcal{N}^{\infty}(T)\leq\mathds{L},$
where the bootstrap constants satisfy
$(\mathcal{I}^{0})^{2}+\mathcal{I}^{0}+1<\min\{\Gamma,\mathds{Y},\mathds{L}\},\qquad\Gamma+\mathds{Y}+\mathds{L}\leq e^{T_{0}/5}.$
Starting from large but finite initial data at $T_{0}$,
$\mathcal{O}(T_{0})+\mathcal{F}(T_{0})+\mathcal{N}^{\infty}(T_{0})\lesssim\mathcal{I}^{0},$
the goal is to show that for all future $T$,
$\mathcal{O}(T)+\mathcal{F}(T)+\mathcal{N}^{\infty}(T)\lesssim\mathcal{I}^{0}+1,$
thereby precluding any geometric blow-up and yielding global-in-time control.
The bootstrap closure relies on a hierarchical chain of control:
$\mathcal{N}^{\infty}\quad\rightarrow\quad\mathcal{F}\quad\rightarrow\quad\mathcal{O}.$
The logic is as follows:
  1. (1)
    From $\mathcal{N}^{\infty}$ to $\mathcal{F}$. The lapse function NN and the deformation tensor $\Sigma$ are controlled point-wise on each slice via elliptic regularity and Sobolev inequalities. $\mathcal{F}$ is controlled by integration the equation for $\Sigma$. See proposition 4.1 and theorem 4.1 in section 4 for a detailed exposition.
  2. (2)
    From $\mathcal{F}$ to $\mathcal{O}$. Once the geometric entropy $\mathcal{F}$ is under control, one derives energy inequalities for the curvature quantities, leading to a uniform bound for $\mathcal{O}$. See the proof of the theorem 4.2 in the section 4.3 for a detailed exposition.
The crucial damping mechanism arises in the mixed term
$\sum_{I\leq 2}e^{-T}\int_{M(T)}N\,\nabla^{I}\mathfrak{T}\,\nabla^{I}\Sigma\,\mu_{g},$
which appears in the $\mathcal{F}$-energy identity (lower order). This term satisfies the bound
$\Big|\sum_{I\leq 2}e^{-T}\!\int_{M(T)}N\,\nabla^{I}\mathfrak{T}\,\nabla^{I}\Sigma\,\mu_{g}\Big|\lesssim e^{-T}\mathcal{O}^{2}.$
The universal damping factor $e^{-T}$ allows one to close the estimate for $\mathcal{F}$:
$\mathcal{F}\lesssim 1+\mathcal{I}_{0}+\mathcal{O}^{2}.$
The appearance of $e^{-T}$ is the essential difference from the $\Lambda=0$ case: when $\Lambda=0$, no such decay is present, forcing $\mathfrak{T}$ to be chosen small to maintain the bootstrap. This recovers the small-data regime treated by [3] (since by the result of Ye [52], smallness of $\mathfrak{T}$ along with the geometric conditions available in this framework would imply the existence of a hyperbolic metric on $M$). In contrast, for $\Lambda>0$, the exponential damping allows one to treat large data, provided the initial time $T_{0}=a\gg 1$.
The importance of $\mathcal{F}$ becomes evident in the highest-order curvature estimates. One encounters Riccati-type nonlinearities of the form
$\sum_{I\leq 3}\sum_{m=0}^{I}\!\!\sum_{J_{1}+J_{2}+J_{3}=m}e^{-T}\!\!\int_{M(T)}\nabla^{J_{1}}N\,\nabla^{J_{2}+1}\Sigma\,\nabla^{J_{3}+I-m}\Sigma\,\nabla^{I+1}\Sigma\,\mu_{g},$
which can be estimated as
$\lesssim e^{-(1+\gamma)T}\mathcal{F}\mathcal{O}^{2}\lesssim e^{-(1+\gamma)T}\mathcal{O}^{4}\lesssim e^{-(1+\gamma)T}\Gamma^{4}.$
Upon integration in time from $T_{0}$ to $T>T_{0}$, the energy contribution is $\lesssim\bigg[e^{-(1+\gamma)T_{0}}-e^{-(1+\gamma)T}\bigg]\Gamma^{4}\lesssim 1$, which closes the $\mathcal{O}$-estimate. In particular, observe that for an arbitrarily large but finite $\mathcal{I}_{0}$, there always exists a large $T_{0}=a$ leading to this estimate since the relative smallness is dictated by $\mathcal{I}^{0}e^{-a/10}<1$ in the main theorem 1.1. The bootstrap assumptions are thereby improved, completing the argument. This explains the idea behind the relative smallness and allowance for large data.
To summarize, the exponential factor $e^{-T}$- a direct consequence of $\Lambda>0$ introduces a universal dissipative mechanism that couples curvature, geometry, and gauge in a self-stabilizing fashion provided that $T_{0}=a\gg 1$. The interaction between $\mathcal{O}$, $\mathcal{F}$, and $\mathcal{N}^{\infty}$ ensures that curvature control enhances geometric regularity, which in turn stabilizes the gauge, feeding back to maintain bounded curvature. Simultaneously, integrating the transport equation
$\frac{\partial g_{ij}}{\partial T}\approx e^{-T}N\Sigma_{ij}-2\!\left(1-\frac{N}{n}\right)g_{ij},$
shows that the metric coefficients evolve smoothly without degenerating Sobolev constants (see propositions 3.13.2, 3.3, and 3.4 in sections 3.4 and 3.5, respectively for detailed proofs). Thus, the combined bootstrap–damping mechanism yields a globally controlled evolution for large initial data in the $\Lambda>0$ regime.
Remark 5.
The analytical framework developed herein extends to the case in which the underlying manifold MM admits a positive Yamabe invariant, i.e., $\sigma(M)>0$, which includes in particular the setting with positive cosmological constant $\Lambda>0$. Under an additional spectral condition on the Laplace–Beltrami operator associated to the evolving metric, ensuring the uniqueness of the solution to the elliptic lapse equation, we establish convergence of the metric to one with constant positive scalar curvature.
This spectral condition, while technical, is natural in view of its role in guaranteeing elliptic solvability and uniqueness. In the absence of such a condition, uniqueness of the limiting metric generally fails. This phenomenon is closely related to the non-uniqueness observed in the convergence behavior of the Yamabe flow on manifolds with positive Yamabe type, as elucidated in the work of Brendle [6], where the flow is shown to converge to a constant scalar curvature metric, but the limit depends nontrivially on the initial data due to the lack of uniqueness in the conformal class.
In our context, the underlying cause of non-uniqueness is distinct and originates from the failure of uniqueness of solutions to the lapse equation, rather than the conformal degeneracy of the space of constant scalar curvature metrics. We emphasize that, even in the presence of convergence, the limiting metric is generally not conformal to the initial metric. For completeness, and to illustrate the applicability of our method in this setting, we state the corresponding convergence result below; the proof proceeds with only minor modifications to the arguments presented in the negative Yamabe case.
First, note that in case of $\sigma(M)>0$, the renormalized lapse equation takes the following form
(30) $-\Delta_{g}(1-\frac{N}{n})+(|\Sigma|^{2}-\frac{1}{n})(1-\frac{N}{n})=-|\Sigma|^{2},~\Delta_{g}:=g^{ij}\nabla_{i}\nabla_{j}.$
A subtle technical issue arises from the analysis of the elliptic equation (30) for the renormalized lapse $1-\frac{N}{n}$. Namely, the associated operator
$\Delta_{g}+n\left(|\Sigma|^{2}-\frac{1}{n}\right)$
may admit a nontrivial finite-dimensional kernel. To proceed, we impose the technical assumption that a solution to (30) exists. Note that such a solution need not be unique: even in the case of de Sitter spacetime, non-uniqueness of the lapse function is known for certain foliations.
Importantly, since the renormalized lapse function $1-\frac{N}{n}$ appears as a source in the evolution equations for $(\Sigma,\mathfrak{T})$, this ambiguity propagates into the solution of the full reduced Einstein system. As a result, the limit geometry determined by the evolution is non-unique, reflecting the intrinsic degeneracy introduced by the kernel of the elliptic operator.
Theorem 1.2 (Semi-global existence for $\Lambda>0$ and $\sigma(M)>0$).
[…]
$\sigma(M)>0.$
Fix $\mathcal{I}^{0}>0$. Then there exists a constant
$a=a(\mathcal{I}^{0})>0$
sufficiently large such that
$\mathcal{I}^{0}e^{-\frac{a}{10}}<1.$
Suppose that the initial data set
$(g_{0},\Sigma_{0})\in C^{\infty}\big(\mathrm{Sym}^{2}(T^{*}M)\big)\times C^{\infty}\big(\mathrm{Sym}^{2}_{0}(T^{*}M)\big)$
for the reduced Einstein-$\Lambda$ system in constant mean curvature (CMC) transported spatial gauge satisfies the constraint equations and further verifies the bounds
(31) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},$
and
(32) $\sum_{0\leq I\leq 3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{0\leq I\leq 2}\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}+\sum_{0\leq I\leq 2}\|e^{a}\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\leq\mathcal{I}^{0},$
where $\xi_{0}$ is a fixed smooth background Riemannian metric on $M$ and $C>0$ is a uniform constant.
Assume further that there exists a solution to the lapse equation corresponding to the positive Yamabe case (see Equation (30)).
Then the maximal developement
$T\mapsto(g(T),\Sigma(T))\in C^{\infty}\big(M;\mathrm{Sym}^{2}(T^{*}M)\big)\times C^{\infty}\big(M;\mathrm{Sym}^{2}_{0}(T^{*}M)\big),$
to the Einstein-$\Lambda$ system with initial data $(g_{0},\Sigma_{0})$ at CMC Newtonian time $T_{0}=a$, exists globally for all
$T\in[T_{0},\infty).$
Moreover, the solution exhibits the following asymptotic behavior as $T\to\infty$:
  1. (i)
    The tensor $\Sigma(T)$ decays to zero in appropriate Sobolev norms;
  2. (ii)
    The metrics $g(T)$ converge smoothly to a limit metric $\widetilde{g}$ on $M$ with constant positive scalar curvature,
    $R(\widetilde{g})(x)=\frac{n-1}{n},\quad\forall x\in M.$
Finally, the limit metric $\widetilde{g}$ is non-unique precisely when the lapse equation admits non-unique solutions.
Conjecture 1.2.
Suppose Consider the following data-set for Einstein-$\Lambda$ vacuum system.
1.7. Consequences for Geometrization
An important conceptual motivation underlying the present work concerns the potential role of the Einstein evolution equations in the geometric classification of closed $3$-manifolds. Specifically, we consider the relevance of the long-time behavior of solutions to the Einstein flow in connection with Thurston’s geometrization conjecture. In this discussion, we assume throughout that the initial data satisfy the regularity assumptions stipulated in the main theorem.
A central open problem in mathematical general relativity is whether the Einstein equations on globally hyperbolic, spatially compact cosmological spacetimes can dynamically implement a form of geometrization of the underlying spatial manifold. In particular, we consider the case in which the Cauchy hypersurface $M$ is a closed, oriented, connected $3$-manifold of negative Yamabe type, i.e., with $\sigma(M)\leq 0$.
The geometrization program in the setting of vacuum cosmological spacetimes was first proposed in the foundational works of Fischer–Moncrief [19, 20], with further developments by Anderson [5], who studied the asymptotic behavior of the vacuum Einstein flow and its possible connection to the Thurston decomposition. In the idealized scenario, one anticipates that the spatial metric induced on constant mean curvature (CMC) hypersurfaces asymptotically converges to a metric structure realizing a geometric decomposition of $M$.
Following Anderson [5], we recall the following definition of geometrization:
Definition 1 (Geometrization).
Let $M$ be a closed, oriented, connected $3$-manifold with $\sigma(M)\leq 0$. A weak geometrization of $M$ is a decomposition
$M=H\cup G,$
where:
  • $H$ is a finite collection of embedded, complete, connected, finite-volume hyperbolic manifolds;
  • $G$ is a finite collection of embedded, connected graph manifolds;
  • the union is along a finite collection of embedded tori $\{T_{i}\}$, with $\partial G=\partial H=\bigsqcup_{i}T_{i}$.
A strong geometrization is a weak geometrization in which each gluing torus $T_{i}$ is incompressible, i.e., the inclusion-induced homomorphism $\pi_{1}(T_{i})\to\pi_{1}(M)$ is injective.
Results due to Ringström [29] show that for a large class of initial data, including arbitrary spatial topologies, the Einstein-$\Lambda$ flow becomes essentially local in nature, and the large-scale topological structure of $M$ becomes invisible to the dynamics. That is, the spatial topology is not imprinted on the asymptotic geometry of the evolving spacetime. We prove indeed this is the case for large initial data.
A heuristic explanation for this phenomenon is furnished by comparison with a classical theorem of Cheeger and Gromov [7, 8], which characterizes volume collapse of Riemannian manifolds with bounded curvature. Specifically, if a sequence $\{(M_{j},g_{j})\}$ of compact Riemannian $3$-manifolds exhibits collapse with two-sided curvature bounds, then $M_{j}$ admits an $\mathcal{F}$-structure of positive rank, and hence is a graph manifold for sufficiently large $j$. Thus, any dynamical mechanism aiming to realize geometrization via Einstein evolution must, at minimum, produce volume collapse with bounded curvature in the graph component $G\subset M$ for large times.
[…]
Corollary 1.2 (No-Collapse Theorem).
Let $B(1)\subset M$ denote a geodesic ball of unit radius (scale $1$ in injectivity radius scale) with respect to the induced metric at CMC time $T\in[T_{0},\infty)$. Then the volume $\operatorname{Vol}(B(1))_{T}$ remains uniformly bounded above by the initial volume $\operatorname{Vol}(B(1))_{T_{0}}$ for all $T\geq T_{0}$.
[…]
The question of whether the Einstein equations with vanishing cosmological constant $\Lambda=0$ can effectuate a geometric decomposition of $M$ remains open and considerably more subtle. Unlike the Ricci flow, where Perelman’s resolution of the geometrization conjecture [40, 41, 42] crucially exploits the parabolic smoothing structure of the flow, the Einstein equations exhibit hyperbolic behavior, and their global-in-time analysis poses formidable technical challenges. Nonetheless, recent advances in the analysis of nonlinear wave equations in the large-data regime, such as [9], offer promising avenues toward addressing these questions, and it is conceivable that similar techniques could be adapted to the cosmological setting.
Ringström [29] provided a detailed local analysis of the Cauchy problem in general relativity with a positive cosmological constant and proposed a formal framework for quantifying the asymptotic inaccessibility of spatial topology by causal observers. His notion supports the view that long-time evolution under the Einstein-$\Lambda$ flow generically suppresses detectable topological information in the causal past of observers.
Definition 2.
[…]
Based on this definition, Ringström proposed the following conjecture:
Conjecture 1.3 ([29]).
[…]
We provide an affirmative to this conjecture in section 5.
1.8. Cosmological Implications
Although our analysis is grounded in the geometric analysis of the Einstein field equations with $\Lambda>0$, we briefly comment on an implication of our results for a longstanding open problem in mathematical cosmology. A central hypothesis in physical cosmology is the so-called cosmological principle, which posits that, at sufficiently large spatial scales, the universe is approximately homogeneous and isotropic. This assumption underlies the derivation of the Friedmann–Lemaître–Robertson–Walker (FLRW) class of solutions, in which constant sectional curvature geometries model the spatial slices.
From a rigorous mathematical perspective, a natural question is whether such large-scale homogeneity and isotropy can arise dynamically from generic initial data under the evolution dictated by the Einstein equations. In particular, one may ask whether spatial slices, even if initially far from being homogeneous or isotropic, can converge in an appropriate topology to a constant sectional curvature geometry—e.g., hyperbolic geometry—along the cosmological constant–driven CMC evolution. This idea was conjectured in [33], where it was suggested that the long-time dynamics of the Einstein-$\Lambda$ flow may select the hyperbolic metric as a universal attractor.
Our analysis, based on Theorem 1.1, provides a rigorous counterexample to this conjecture. Indeed, we work with the most general class of closed, connected, oriented three-manifolds of negative Yamabe type, encompassing all topological classes listed in Cases 1–3 of section 1. For such generic topologies, we demonstrate that no form of geometric convergence of the rescaled spatial metric to a hyperbolic (i.e., maximally symmetric) geometry occurs.
Corollary 1.3 (Failure of the Cosmological Principle as a Derived Consequence).
[…]
A proof of this statement is immediate from the structure of the limiting data described in Remark 12.
[…]
Acknowledgement
I thank Professor S-T Yau for many discussions regarding the scalar curvature geometry and Professor V. Moncrief for numerous discussions regarding Einstein flow. I thank the reviewer for the meticulous review, which substantially improved the quality of the article. This work was supported by the Center of Mathematical Sciences and Applications, Department of Mathematics at Harvard University and by BIMSA of YMSC at Tsinghua University.
2. Preliminaries
[…]
$\widetilde{M}\cong\mathbb{R}\times M,$
where each leaf $\{t\}\times M$ is an orientable, compact $n$-dimensional Cauchy hypersurface diffeomorphic to $M$. The Lorentzian metric $\widehat{g}$ induces a Riemannian metric $g(t)$ on each slice $M(t):=\{t\}\times M$.
The ADM decomposition expresses the future-directed timelike unit normal vector field $\widehat{\mathsfbf{n}}$ orthogonal to the slices and the vector field $\partial_{t}$ generating the foliation as
$\partial_{t}=N\widehat{\mathsfbf{n}}+X,$
where $N:\widetilde{M}\to(0,\infty)$ is the lapse function, and $X\in\Gamma(TM)$ is the shift vector field, both of which possess the requisite regularity dictated by the function spaces to be specified later. The Lorentzian metric in the adapted coordinate system $\{x^{\alpha}\}_{\alpha=0}^{n}=\{t,x^{i}\}_{i=1}^{n}$ admits the canonical form
(33) $\widehat{g}=-N^{2}dt\otimes dt+g_{ij}(dx^{i}+X^{i}dt)\otimes(dx^{j}+X^{j}dt),$
where $g_{ij}(t,x)$ is the induced Riemannian metric on $M(t)$.
To quantify the extrinsic geometry of the embedding $M(t)\hookrightarrow\widetilde{M}$, we define the second fundamental form $k\in\Gamma(S^{2}T^{*}M)$ by
$k_{ij}:=-\langle\nabla_{\partial_{i}}\widehat{\mathsfbf{n}},\partial_{j}\rangle=-\frac{1}{2N}\bigl(\partial_{t}g_{ij}-\mathscr{L}_{X}g_{ij}\bigr),$
where $\nabla$ denotes the Levi-Civita connection of $\widehat{g}$, and $\mathscr{L}_{X}$ is the Lie derivative along $X$. The mean extrinsic curvature, or trace of $k$, is given by
$\tau:=\mathrm{tr}_{g}k=g^{ij}k_{ij},$
with $g^{ij}$ denoting the inverse metric on $M(t)$.
The vacuum Einstein field equations with cosmological constant $\Lambda\in\mathbb{R}$,
(34) $\mathrm{Ric}(\widehat{g})-\frac{1}{2}R(\widehat{g})\widehat{g}+\Lambda\widehat{g}=0,$
admit an equivalent formulation in terms of the Cauchy data $(g,k)$ on each hypersurface $M(t)$ as the coupled evolution and constraint system:
(35) $\partial_{t}g_{ij}$ $=-2Nk_{ij}+\mathscr{L}_{X}g_{ij},$
(36) $\partial_{t}k_{ij}$ $=-\nabla_{i}\nabla_{j}N+N\left(R_{ij}+\tau k_{ij}-2k_{i}^{m}k_{mj}-\frac{2\Lambda}{n-1}g_{ij}\right)+\mathscr{L}_{X}k_{ij},$
(37) $R-|k|_{g}^{2}+\tau^{2}$ $=2\Lambda,$
(38) $\nabla^{i}k_{ij}-\nabla_{j}\tau$ $=0,$
where $\nabla$ denotes the Levi-Civita connection of $g$, $R_{ij}$ and $R$ are the Ricci and scalar curvatures of $g$, respectively, and $|k|_{g}^{2}=g^{im}g^{jn}k_{ij}k_{mn}$. Here and henceforth, indices are raised and lowered by $g$ and its inverse.
We denote by $\Gamma[g]$ the Christoffel symbols associated to $g$, and by $\mathrm{Riem}$ and $\mathrm{Ric}$ the corresponding Riemann and Ricci curvature tensors on $M$. The (negative-semidefinite) Laplace–Beltrami operator acting on functions $f\in C^{\infty}(M)$ is given by
$\Delta_{g}f:=g^{ij}\nabla_{i}\nabla_{j}f,$
where the sign convention ensures that the spectrum of $\Delta_{g}$ is nonnegative.
A central object in our analysis is the Lichnerowicz Laplacian
$\mathcal{L}_{g}:\Gamma(S^{2}T^{*}M)\to\Gamma(S^{2}T^{*}M),$
defined by
$\mathcal{L}_{g}u_{ij}:=-\Delta_{g}u_{ij}-2R_{i}{}^{k}{}_{j}{}^{l}u_{kl},$
where the curvature operator acts via the Riemann curvature tensor. For metrics $g$ which are stable Einstein metrics, the operator $\mathcal{L}_{g}$ is elliptic and possesses strictly positive spectrum on the space of symmetric 2-tensors modulo gauge.
Finally, for nonnegative scalar functions $f,h:[0,\infty)\to\mathbb{R}_{\geq 0}$, we write $f(t)\lesssim h(t)$ to signify the existence of a constant $C>0$, independent of $t$ but possibly dependent on the fixed background geometry and dimension $n$, such that
$f(t)\leq Ch(t),$
and $f(t)\approx h(t)$ if there exist constants $0<C_{1}\leq C_{2}<\infty$ such that
$C_{1}h(t)\leq f(t)\leq C_{2}h(t).$
Throughout, the spaces of smooth symmetric covariant 2-tensors and vector fields on $M$ are denoted by $S^{0}_{2}(M)$ and $\mathfrak{X}(M)$, respectively. The trace-free and transverse-traceless parts of a symmetric 2-tensor $A$ are denoted by $\widehat{A}$ and $A^{TT}$, respectively. We will also frequently denote the Cauchy slice at constant mean curvature time $T$ by $M(T)$. In addition, we frequently use the following transport identity which is a consequence of the transport theorem [32]
(39) $\frac{d}{dT}\int_{M(T)}f\mu_{g}=\int_{M(T)}(\partial_{T}f+\frac{1}{2}f\mbox{tr}_{g}\partial_{T}g)\mu_{g}.$
3. Gauge fixed Einstein’s equations
We study the Cauchy problem in constant mean extrinsic curvature (CMC) gauge. To this end, we assume that the Lorentzian spacetime diffeomorphic to $M\times\mathbb{R}$ admits a constant mean curvature Cauchy slice.
Now recall the vacuum Einstein’s field equations with a cosmological constant $\Lambda$
(40) $\text{Ric}(\widehat{g})_{\mu\nu}-\frac{1}{2}R(\widehat{g})+\Lambda\widehat{g}_{\mu\nu}=0$
expressed in $3+1$ formulation in a local chart $(t,x^{1},x^{2},x^{3})$
(41) $\partial_{t}g_{ij}=-2Nk_{ij}+(\mathscr{L}_{X}g)_{ij},$
(42) $\partial_{t}k_{ij}=-\nabla_{i}\nabla_{j}N+N(R_{ij}+\mbox{tr}_{g}kk_{ij}-2k^{k}_{i}k_{jk}-\frac{2\Lambda}{n-1}g_{ij})+(\mathscr{L}_{X}k)_{ij},$
(43) $2\Lambda=R(g)-|k|^{2}_{g}+(\mbox{tr}_{g}k)^{2},$
(44) $0=\nabla^{i}k_{ij}-\nabla_{j}\mbox{tr}_{g}k.$
This evolution-constraint system leads to the following initial value problem
Definition 3.
Initial data for 40 consist of an n dimensional manifold $M$, a Riemannian metric $g$, a covariant 2-tensor $k$, all assumed to be smooth and to satisfy
(45) $2\Lambda=R(g)-|k|^{2}_{g}+(\mbox{tr}_{g}k)^{2},$
(46) $0=\nabla^{i}k_{ij}-\nabla_{j}\mbox{tr}_{g}k.$
[…]
Now we define the CMC gauge. First, recall that
Definition 4 (Constant Mean Curvature (CMC) Gauge).
[…]
$\mathrm{tr}_{g_{t}}k_{t}=\tau(t)=t,$
where $g_{t}$ denotes the induced Riemannian metric on $M_{t}$. In this gauge, the time coordinate function $t$ is chosen such that it coincides with the mean curvature of the hypersurface $M_{t}$.
The definition automatically requires that the mean extrinsic curvature of level sets of the time be constant on the level sets i.e., $\nabla\tau=0$. More precisely the mean extrinsic curvature $\tau=\mbox{tr}_{g}k$ of $M$ verifies the following
(47) $\tau=\mbox{tr}_{g}k=\text{monotonic function of $t$ alone}.$
In particular, we choose
(48) $t=\tau.$
The first analytical issue to address is the existence of a constant mean curvature (CMC) hypersurface $M$ within the spacetime manifold $M\times\mathbb{R}$. Unlike the setting of stability problems, where one typically begins with a known background solution possessing CMC slices, often of negative Einstein type, our situation lacks such an a priori reference geometry. In stability analyses, the presence of a negative Einstein background ensures the existence of CMC slices via a rigidity argument, whereby small perturbations of the data continue to admit CMC foliations. However, in the present context, no such background solution is assumed to be available. Rather, the goal is to determine whether such solutions exist at all, whether generic solutions asymptotically approach them, and whether any geometric or analytic obstructions prevent this convergence.
Consequently, we impose as a working hypothesis that the spacetime admits at least one CMC hypersurface. This assumption is motivated by the asymptotic structure of the models under consideration, which exhibit a big-bang-type singularity in the past. Specifically, we assume that past-directed inextendible causal geodesics terminate in a curvature singularity, in the sense of the $C^{\infty}$ formulation of the strong cosmic censorship conjecture in the vacuum cosmological setting. Under these conditions, the following result of Marsden and Tipler ensures the existence of a CMC hypersurface:
Theorem ([31], Theorem 6).
[…]
In the CMC gauge, the first notice a few basic important points.
Claim 1.
[…]
Proof.
The proof proceeds by a contradiction argument based on the Hamiltonian constraint and the structure of the space of initial data. We aim to establish the impossibility of the existence of a maximal slice (i.e., one for which $\tau=0$) in the future development of an initially expanding solution.
We begin by recalling that any symmetric $(0,2)$-tensor $k$ on $M$ can be orthogonally decomposed with respect to the standard York splitting (cf. [26]):
(49) $k=k^{\mathrm{TT}}+\frac{\tau}{3}g+\left(\mathscr{L}_{Z}g-\frac{2}{3}(\operatorname{div}_{g}Z)g\right),$
where $k^{TT}$ denotes the transverse-traceless (TT) part of $k$, satisfying $tr_{g}k^{TT}$ and $\mbox{div }_{g}k^{TT}=0$, and $Z$ is a smooth vector field on $M$.
Suppose $k$ satisfies the momentum constraint:
(50) $\nabla^{j}k_{ij}=0.$
We show that the non-TT part involving the vector field Z must vanish. To that end, consider the $L^{2}$-inner product of the momentum constraint with the vector field Z, and apply Stokes’ theorem:
$0$ $=\int_{M}\left(Z^{i}\nabla^{j}k_{ij}+Z^{j}\nabla^{i}k_{ji}\right)\mu_{g}\$$=\int_{M}(\nabla^{i}Z^{j}+\nabla^{j}Z^{i})\left(\nabla_{i}Z_{j}+\nabla_{j}Z_{i}-\frac{2}{3}(\operatorname{div}gZ)g{ij}\right)\mu_{g}$
$=\int_{M}\left|\mathscr{L}_{Z}g-\frac{2}{3}(\operatorname{div}_{g}Z)g\right|^{2}\mu_{g}.$
Since the integrand is non-negative, this integral vanishes if and only if
(51) $\mathscr{L}_{Z}g-\frac{2}{3}(\operatorname{div}_{g}Z)g=0.$
Thus, the extrinsic curvature simplifies to:
(52) $k=k^{\mathrm{TT}}+\frac{\tau}{n}g.$
We now invoke the Hamiltonian constraint on a CMC hypersurface $M_{t}$, which reads:
(53) $R(g)-|k|^{2}+\tau^{2}$ $=2\Lambda\ \Rightarrow R(g)-\left|k^{\mathrm{TT}}+\frac{\tau}{n}g\right|^{2}+\tau^{2}=2\Lambda.$
Expanding the norm
(54) $|k^{\mathrm{TT}}+\frac{\tau}{n}g|^{2}=|k^{\mathrm{TT}}|^{2}+\frac{2\tau}{n}\operatorname{tr}_{g}k^{\mathrm{TT}}+\frac{\tau^{2}}{n^{2}}\operatorname{tr}_{g}g=|k^{\mathrm{TT}}|^{2}+\frac{\tau^{2}}{n},$
and using $\mbox{tr}_{g}k^{TT}=0$ and $\mbox{tr}_{g}g=n$, we obtain:
(55) $R(g)-|k^{\mathrm{TT}}|^{2}=2\Lambda-\tau^{2}\left(1-\frac{1}{n}\right).$
For $n=3$, this simplifies to:
(56) $R(g)-|k^{\mathrm{TT}}|^{2}=2\Lambda-\frac{2}{3}\tau^{2}.$
Now, the hypothesis that $M$ is of negative Yamabe type implies that for all conformal representatives $g$ in the conformal class $[g]$, the scalar curvature $R(g)$ cannot be everywhere nonnegative. In particular, for all metrics $g$ conformally related to a constant negative scalar curvature metric, $R(g)<0$ somewhere. Thus, the left-hand side is strictly negative at some point, which enforces the inequality:
(57) $\quad\tau^{2}>3\Lambda.$
Since $\tau$ is constant on each CMC hypersurface, this implies that $\tau^{2}>3\Lambda$, hence, in particular, $\tau$ cannot vanish. Therefore, a maximal slice (where $\tau=0$) cannot form during the evolution. Finally, since $\tau(t)$ is monotonically increasing in time by the choice of CMC gauge (cf. Equation (47)), any initially expanding solution (i.e., with $\tau<0$) must remain strictly expanding throughout its domain of existence, with $\tau<0$ for all $t$. This concludes the proof. ∎
Remark 6.
For convenience, one could simply set $\Lambda=n(n+1)$.
A central function in the construction of a dimensionless dynamical formulation of the Einstein vacuum equations with cosmological constant $\Lambda$ is the use of a natural geometric scaling governed by the mean curvature. Consider the entity
$\varphi^{2}:=\tau^{2}-3\Lambda=\tau^{2}-3n(n+1)>0,$
which is strictly positive and decays monotonically toward the future in a chosen temporal gauge (claim 1). The function $\varphi:=\sqrt{\tau^{2}-3n(n+1)}$ serves as a natural scaling factor for the spatial geometry and lapse-shift data. To obtain a dimensionally consistent and dynamically well-posed evolution system, one must perform a rescaling of the Einstein field equations in terms of $\varphi$.
We adopt the dimensional convention of [3], wherein the local spatial coordinates $\{x^{i}\}$ on the Cauchy hypersurface $M$ are taken to be dimensionless. Under this convention, the geometric and gauge quantities possess the following physical dimensions:
(58) $[g_{ij}]=L^{2},\quad[k_{ij}]=L,\quad[X^{i}]=L,\quad[N]=L^{2},\quad[\tau]=L^{-1},\quad[\varphi^{2}]=L^{-2},$
where $L$ denotes the dimension of length.
To facilitate the introduction of dimensionless variables, we adopt the notational convention that dimensional quantities are denoted with a tilde, while their dimensionless counterparts are written without decoration. The natural rescaling is then given by
(59) $\widetilde{g}_{ij}=\frac{1}{\varphi^{2}}g_{ij},\qquad\widetilde{N}=\frac{1}{\varphi^{2}}N,\qquad\widetilde{X}^{i}=\frac{1}{\varphi}X^{i},\qquad\widetilde{k}_{ij}^{TT}=\frac{1}{\varphi}\Sigma_{ij},$
where $\Sigma_{ij}$ denotes the trace-free part of the extrinsic curvature, and we define $\varphi=-\sqrt{\tau^{2}-3n(n+1)}$ so that $\varphi/\tau>0$.
We now introduce a reparametrization of time via a new variable $T=T(\tau)$, defined implicitly by the vector field transformation
(60) $\partial_{T}=-\frac{\varphi^{2}}{\tau}\partial_{\tau}.$
Integrating this relation yields
(61) $\varphi(T)=\varphi(T_{0})e^{T_{0}}e^{-T},$
for some fixed initial time $T_{0}$. By setting the normalization condition $\varphi(T_{0})e^{T_{0}}=-C$ for a constant $C$ of dimension $L^{-1}$, we obtain explicit expressions for $\varphi$ and $\tau$ as functions of the new time coordinate:
(62) $\varphi(T)=-Ce^{-T},\qquad\tau(T)=-\sqrt{C^{2}e^{-2T}+3n(n+1)}.$
The new time variable $T$ ranges over the entire real line, i.e., $T\in(-\infty,\infty)$, and behaves analogously to Newtonian time in the rescaled formulation.
Moreover, we record the asymptotic estimate
(63) $\frac{\varphi(T)}{\tau(T)}=\frac{Ce^{-T}}{\sqrt{C^{2}e^{-2T}+3n(n+1)}}\leq\frac{Ce^{-T}}{\sqrt{3n(n+1)}}\sim n^{-\frac{1}{2}}(n+1)^{-\frac{1}{2}}e^{-T},$
demonstrating the exponential decay of the ratio $\varphi/\tau$ in the future time direction.
The rescaled Einstein’s equations in the CMC gauge can be cast into the following system
(64) $\frac{\partial g_{ij}}{\partial T}=-\frac{2\varphi}{\tau}N\Sigma_{ij}-2(1-\frac{N}{n})g_{ij}-\frac{\varphi}{\tau}(\mathscr{L}_{X}g)_{ij}$
(65) $\frac{\partial\Sigma_{ij}}{\partial T}=-(n-1)\Sigma_{ij}-\frac{\varphi}{\tau}N(\text{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij})+\frac{\varphi}{\tau}\nabla_{i}\nabla_{j}(\frac{N}{n}-1)+\frac{2\varphi}{\tau}N\Sigma_{ik}\Sigma^{k}_{j}$
(66) $-\frac{\varphi}{n\tau}(\frac{N}{n}-1)g_{ij}-(n-2)(\frac{N}{n}-1)\Sigma_{ij}-\frac{\varphi}{\tau}(\mathscr{L}_{X}\Sigma)_{ij}$
(67) $R+\frac{n-1}{n}-|\Sigma|^{2}=0,$
(68) $\nabla_{j}\Sigma^{j}_{i}=0.$
For the purpose of our study, this system is not suitable. Instead of treating this system as a coupled weakly wave system (which would inevitably require invoking spatial harmonic gauge to turn the system into a strong hyperbolic system), we will treat the equation for the metric components $g_{ij}$ as transport equations. We define the following new entity that we shall study
(69) $\mathfrak{T}_{ij}:=\text{Ric}_{ij}-\frac{1}{n}R(g)g_{ij}.$
We obtain the following manifestly wave system for $(\Sigma,\mathfrak{T})$.
Remark 7.
We note that Choquet-Bruhat and York [1] considered a similar coupled system for $\Sigma$ and Ric.
Lemma 3.1 (Coupled Wave System for $(\Sigma,\mathfrak{T})$).
Let $(g,\Sigma)$ satisfy the reduced Einstein evolution equations (64)–(68) in constant mean curvature (CMC) and transported spatial gauge. Then the pair $(\Sigma,\mathfrak{T})$, consisting respectively of the transverse-traceless part of the second fundamental form and the obstruction tensor $\mathfrak{T}$, satisfies the following system of coupled second-order quasilinear hyperbolic equations, modulo a spatial diffeomorphism $\Psi$ generated by a smooth shift vector field $X$.
The evolution equations for $\Sigma_{ij}$ and $\mathfrak{T}_{ij}$ read:
$\partial_{T}\Sigma_{ij}$ $=-\frac{\varphi}{\tau}(\mathscr{L}_{X}\Sigma)_{ij}-(n-1)\Sigma_{ij}-\frac{\varphi}{\tau}N\mathfrak{T}_{ij}+\frac{\varphi}{\tau}\nabla_{i}\nabla_{j}\left(\frac{N}{n}-1\right)+\frac{2\varphi}{\tau}N\Sigma_{ik}\Sigma^{k}{}_{j}$
(70) $\quad-\frac{\varphi}{n\tau}\left(\frac{N}{n}-1\right)g_{ij}-(n-2)\left(\frac{N}{n}-1\right)\Sigma_{ij},$
[…]
$\quad+n\frac{\varphi}{\tau}\nabla^{l}\nabla_{i}\left(\frac{N}{n}-1\right)\Sigma_{jl}+n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{i}\Sigma_{jl}$
$\quad+n\frac{\varphi}{\tau}\nabla^{l}\nabla_{j}\left(\frac{N}{n}-1\right)\Sigma_{il}+n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{j}\Sigma_{il}$
$\quad-n\frac{\varphi}{\tau}\Delta_{g}\left(\frac{N}{n}-1\right)\Sigma_{ij}-2n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{l}\Sigma_{ij}$
(71) $\quad-\Delta_{g}\left(\frac{N}{n}-1\right)g_{ij}+N\frac{\varphi}{\tau}\left(\mathfrak{T}_{ki}\Sigma^{k}{}_{j}+\mathfrak{T}_{kj}\Sigma^{k}{}_{i}\right)+\frac{2(n-1)}{n^{2}}\left(\frac{N}{n}-1\right)g_{ij}.$
Proof.
The proof follows directly from Einstein’s equations and the variation formula for Ricci and scalar curvature.
There are advantages of working with the pair $(\Sigma,\mathfrak{T})$ instead of $(g,\Sigma)$. Firstly, the equations are manifestly hyperbolic up to a spatial diffeomorphism and one does not need to deal with the Gribov ambiguities associated with the spatial harmonic gauge often used. In particular, the spatial harmonic gauge is not suited for large data problems in the current context without substantial technical machinery.
In the previous section, we have chosen the CMC gauge as the time gauge. We need to choose a spatial gauge to analyze the evolution system 3.1. We choose CMC transported spatial gauge (previously [17, 45] used this gauge to study the dynamical stability of big-bang singularity formation)
3.1. Transported Spatial Coordinates and the Re-Scaled Einstein System
In this section, we introduce the notion of a transportedspatial gauge, which serves as a coordinate choice adapted to the constant mean curvature (CMC) foliation.
Definition 5 (Transported Spatial Gauge).
Let $(x^{1}_{0},x^{2}_{0},x^{3}_{0})$ be local coordinates on an open neighborhood $\Omega\subset M$ of the initial Cauchy hypersurface $\{T^{\prime}=T_{0}\}\subset\widetilde{M}:=\mathbb{R}\times M$. We say that a coordinate system $(T^{\prime},x^{1},x^{2},x^{3})$ on $T_{0},T]\times\Omega$ defines a transported spatial gauge if the spatial coordinate functions are Lie transported along the future-directed unit normal vector field $\widehat{\mathsfbf{n}}$ to the CMC hypersurfaces, i.e.,
(72) $\mathscr{L}_{\widehat{\mathsfbf{n}}}(x^{i})=0,\qquad i=1,2,3,$
with initial condition $(x^{1},x^{2},x^{3})|_{T^{\prime}=T_{0}}=(x^{1}_{0},x^{2}_{0},x^{3}_{0})$.
From (72), and the fact that $\widehat{\mathsfbf{n}}=\frac{1}{N}(\partial_{T^{\prime}}-X^{i}\partial_{i})$ in general coordinates, we deduce:
(73) $\mathscr{L}_{\widehat{\mathsfbf{n}}}(x^{i})=\widehat{\mathsfbf{n}}(x^{i})=\frac{1}{N}(\partial_{T^{\prime}}-X^{j}\partial_{j})(x^{i})=-\frac{1}{N}X^{i}\Rightarrow X^{i}=0.$
Thus, the shift vector field vanishes identically in these coordinates.
In this gauge, the rescaled Einstein evolution equations take a particularly tractable form. Let $\Sigma_{ij}$ denote the trace-free part of the second fundamental form (the shear), and let $\mathfrak{T}_{ij}$ denote the trace-free part of the rescaled Ricci curvature. The evolution equations then read:
(74) $\partial_{T}\Sigma_{ij}$ $=-(n-1)\Sigma_{ij}-\frac{\varphi}{\tau}N\mathfrak{T}_{ij}+\frac{\varphi}{\tau}\nabla_{i}\nabla_{j}\left(\frac{N}{n}-1\right)+\frac{2\varphi}{\tau}N\Sigma_{ik}\Sigma^{k}{}_{j}$
$\quad-\frac{\varphi}{n\tau}\left(\frac{N}{n}-1\right)g_{ij}-(n-2)\left(\frac{N}{n}-1\right)\Sigma_{ij},$
(75) $\partial_{T}\mathfrak{T}_{ij}$ $=-N\frac{\varphi}{\tau}\left(\Delta_{g}\Sigma_{ij}-R^{m}{}_{jli}\Sigma^{l}{}_{m}-R^{m}{}_{ilj}\Sigma^{l}{}_{m}\right)+n\frac{\varphi}{\tau}\nabla^{l}\nabla_{i}\left(\frac{N}{n}-1\right)\Sigma_{jl}$
$\quad+n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{i}\Sigma_{jl}+n\frac{\varphi}{\tau}\nabla^{l}\nabla_{j}\left(\frac{N}{n}-1\right)\Sigma_{il}$
$\quad+n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{j}\Sigma_{il}-n\frac{\varphi}{\tau}\Delta_{g}\left(\frac{N}{n}-1\right)\Sigma_{ij}$
$\quad-2n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{l}\Sigma_{ij}-\Delta_{g}\left(\frac{N}{n}-1\right)g_{ij}$
$\quad+N\frac{\varphi}{\tau}\left(\mathfrak{T}_{ki}\Sigma^{k}{}_{j}+\mathfrak{T}_{kj}\Sigma^{k}{}_{i}\right)+\frac{2(n-1)}{n^{2}}\left(\frac{N}{n}-1\right)g_{ij},$
These equations are supplemented by the elliptic lapse equation:
(76) $-\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|^{2}+\frac{1}{n}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|^{2}.$
3.2. Integration and Norms
Let $(M,g)$ be a smooth, closed, oriented Riemannian $3$-manifold. Fix a smooth partition of unity $\{\xi_{U}\}_{U\in\mathcal{U}}$ subordinate to a finite atlas $\{(U,\varphi_{U})\}_{U\in\mathcal{U}}$ of coordinate charts $\varphi_{U}:U\to\mathbb{R}^{3}$. For a measurable function $f:M\to\mathbb{R}$, the integral of $f$ with respect to the Riemannian volume form $\mu_{g}$ is defined by
(77) $\int_{M}f\,\mu_{g}:=\sum_{U\in\mathcal{U}}\int_{\varphi_{U}(U)}f\circ\varphi_{U}^{-1}(x)\,\xi_{U}\circ\varphi_{U}^{-1}(x)\sqrt{\det g_{ij}(x)}\,dx^{1}dx^{2}dx^{3},$
where $g_{ij}$ denotes the components of the metric in the chart $\varphi_{U}$, and $\det g_{ij}$ is the determinant of the metric matrix.
Given a smooth tensor field $\psi\in\Gamma\left((T^{\otimes L}M)\otimes(T^{*}M)^{\otimes K}\right)$, where $K+L\geq 1$, we define its pointwise norm using the Riemannian metric $g$ by
(78) $\langle\psi(x),\psi(x)\rangle_{g}:=g^{i_{1}j_{1}}\cdots g^{i_{L}j_{L}}g_{m_{1}n_{1}}\cdots g_{m_{K}n_{K}}\,\psi^{m_{1}\dots m_{K}}_{i_{1}\dots i_{L}}(x)\,\psi^{n_{1}\dots n_{K}}_{j_{1}\dots j_{L}}(x).$
This inner product is well-defined and independent of the choice of coordinates. The $L^{p}$ norm of $\psi$ over $(M,g)$ is then defined by
(79) $\|\psi\|_{L^{p}(M)}^{p}:=\int_{M}\left(\langle\psi,\psi\rangle_{g}\right)^{p/2}\,\mu_{g},\qquad 1\leq p<\infty,$
and the essential supremum (or $L^{\infty}$) norm is given by
(80) $\|\psi\|_{L^{\infty}(M)}:=\sup_{x\in M}\left(\langle\psi(x),\psi(x)\rangle_{g}\right)^{1/2}.$
[…]
(81) $\mathcal{O}_{I}:=\begin{cases}\|\nabla^{I}\Sigma\|_{L^{2}(M)}+\|\nabla^{I-1}\mathfrak{T}\|_{L^{2}(M)},&\text{if }I\geq 1,\\ \|\Sigma\|_{L^{2}(M)},&\text{if }I=0.\end{cases}$
We also introduce exponentially weighted norms in time to account for the expanding geometry in cosmological spacetimes. Let $T$ denote the CMC time parameter, and let $\gamma>0$ be a scaling parameter (which will ultimately be chosen to be $\gamma=1$). For $I\leq 2$, define the weighted Sobolev norms
(82) $\mathcal{F}_{I}:=\|e^{\gamma T}\nabla^{I}\Sigma\|_{L^{2}(M)}.$
Define also the pointwise weighted norm:
(83) $\mathcal{N}^{\infty}:=\|e^{\gamma T}\Sigma\|_{L^{\infty}(M)}+\left\|e^{2\gamma T}\left(\frac{N}{n}-1\right)\right\|_{L^{\infty}(M)},$
where $N$ denotes the spacetime lapse function.
The total energy norms used throughout this work are defined by
(84) $\mathcal{O}:=\sum_{I=0}^{3}\mathcal{O}_{I},\qquad\mathcal{F}:=\sum_{I=0}^{2}\mathcal{F}_{I}.$
As will be demonstrated in the analysis below, the optimal exponential weight corresponds to $\gamma=1$, which balances the natural scaling of the lapse and the shear in expanding CMC spacetimes.
3.3. Setting up the bootstrap argument
In order to establish global-in-time control of the solutions of the Einstein vacuum equations with cosmological constant $\Lambda>0$—in the rescaled variables introduced earlier—we employ a bootstrap argument. Specifically, we aim to derive uniform bounds on three scale-invariant norms:
$\mathcal{O},\mathcal{F},\mathcal{N}^{\infty},$
which respectively control curvature energies, renormalized entropy for the transverse-traceless ($TT$) second fundamental form, and point-wise norm of gauge variable and $TT$ tensor $\Sigma$. These norms are defined in the rescaled setting associated with the dynamical time variable $T\in[T_{0},T_{\infty})$.
We assume the initial data at time $T=T_{0}$ satisfies the bound
(85) $\mathcal{O}(T_{0})+\mathcal{F}(T_{0})+\mathcal{N}^{\infty}(T_{0})\lesssim\mathcal{I}^{0},$
where $\mathcal{I}^{0}\gg 1$ denotes the (large) initial data size. Our main objective is to prove that for all $T\in[T_{0},T_{\infty}]$, the combined norm satisfies the uniform a priori estimate
(86) $\mathcal{O}(T)+\mathcal{F}(T)+\mathcal{N}^{\infty}(T)\lesssim\mathcal{I}^{0}+1.$
This suffices to propagate the solution globally in time via standard local well-posedness and continuation arguments (see Section 3.1).
To achieve (86), we proceed via a bootstrap scheme. We assume that on a given time interval $T_{0},T]\subset[T_{0},T_{\infty})$, the following bounds hold:
(87) $\mathcal{O}(T)\leq\Gamma,\qquad\mathcal{F}(T)\leq\mathds{Y},\qquad\mathcal{N}^{\infty}(T)\leq\mathds{L},$
where the constants $\Gamma,\mathds{Y},\mathds{L}$ satisfy
(88) $(\mathcal{I}^{0})^{2}+\mathcal{I}^{0}+1<\min\{\Gamma,\mathds{Y},\mathds{L}\},$
and are additionally constrained by
(89) $\Gamma+\mathds{Y}+\mathds{L}\leq e^{\frac{T_{0}}{10}}.$
Define the set of admissible bootstrap times:
[…]
Our aim is to prove that $\mathcal{S}=[T_{0},T_{\infty}]$ by showing that $\mathcal{S}$ is both open and closed.
Step 1: Non-emptiness
[…]
(90) $\mathcal{O}(T)+\mathcal{F}(T)+\mathcal{N}^{\infty}(T)\lesssim 2\mathcal{I}^{0},\qquad\forall T\in[T_{0},T_{0}+\varepsilon],$
thus implying $\mathcal{S}\neq\{\emptyset\}$.
Step 2: Closedness
The closedness of $\mathcal{S}$ follows from standard continuation results and the fact that the norms $\mathcal{O},\mathcal{F},\mathcal{N}^{\infty}$ are continuous in time. More precisely, the uniform bounds implied by the bootstrap assumptions extend continuously to the supremum of $\mathcal{S}$, which must then also belong to $\mathcal{S}$.
Step 3: Openness
The core of the argument lies in establishing a hierarchy of estimates:
$\mathcal{N}^{\infty}\longrightarrow\mathcal{F}\longrightarrow\mathcal{O},$
that enable one to strictly improve the assumptions in (87). These estimates are derived as follows:
  • The gauge quantities in $\mathcal{N}^{\infty}$ are controlled via elliptic regularity theory and Sobolev inequalities on the spatial slices $\Sigma_{T}$, leading to the estimate
    (91) $\mathcal{N}^{\infty}(T)\lesssim\mathcal{O}(T)+\mathcal{F}(T)^{2}+\mathcal{F}(T).$
  • The norm $\mathcal{F}$, which captures the trace-free part of the second fundamental form, is controlled by integrating the transport equation for $\Sigma$ forward in $T$, yielding
    (92) $\mathcal{F}(T)\lesssim\mathcal{O}(T)+\mathcal{I}^{0}+1.$
  • Finally, $\mathcal{O}$ is estimated via energy methods, applied to the hyperbolic system, giving the uniform bound
    (93) $\mathcal{O}(T)\lesssim\mathcal{I}^{0}+1.$
Combining (91)–(93), and choosing the bootstrap constants sufficiently large relative to $\mathcal{I}^{0}$ yet satisfying (89), we obtain strict improvements of the assumptions in (87). Therefore, by the local well-posedness theory, the solution to the Einsten-$\Lambda$ equation can be extended a bit towards a larger value of $T$, implying the openness of $\mathcal{S}$.
Since $\mathcal{S}\subset[T_{0},T_{\infty}]$ is non-empty, open, and closed, and $T_{0},T_{\infty}]$ is connected, it follows that
$\mathcal{S}=[T_{0},T_{\infty}].$
Thus, the solution exists globally in time, and the norms $\mathcal{O},\mathcal{F},\mathcal{N}^{\infty}$ obey uniform bounds independent of $T_{\infty}$. Sending $T_{\infty}\to\infty$ yields global existence and uniform control for all future times.
3.4. Estimates for the metric components
Proposition 3.1 (Uniform Control of the Metric Volume Density).
Let $(\widetilde{M},\widehat{g}(T))$ be a smooth globally hyperbolic spacetime foliated by constant mean curvature (CMC) slices, and suppose that the metric $g(T)$ evolves according to the Einstein vacuum equations with cosmological constant in the CMC-spatially transported gauge. Let the initial data at $T=T_{0}$ satisfy the assumptions of Theorem 1.1, and assume the bootstrap condition
$\left\|\frac{N}{n}-1\right\|_{L^{\infty}(M)}\leq\mathds{L}e^{-2\gamma T_{0}},\qquad\text{for all }T\in[T_{0},T],$
where $\mathds{L},\gamma>0$ are constants. Then the Riemannian volume density $\mu_{g}:=\sqrt{\det g_{ij}}$ associated with the induced metric $g_{ij}(T)$ satisfies the uniform bounds
$C_{1}\leq\det(g_{ij}(T))\leq C_{2},\qquad\text{for all }T\in[T_{0},T],$
for some positive constants $C_{1},C_{2}>0$ depending only on the initial data at $T_{0}$.
Proof.
We begin with the evolution equation for the spatial metric $g_{ij}$ in CMC-spatially transported gauge:
$\partial_{T}g_{ij}=-\frac{2\varphi}{\tau}N\Sigma_{ij}-2\left(1-\frac{N}{n}\right)g_{ij}.$
Taking the trace with respect to $g^{ij}$, and using the fact that $\operatorname{tr}_{g}\Sigma=0$, we obtain:
$\operatorname{tr}_{g}(\partial_{T}g)=-2n\left(1-\frac{N}{n}\right).$
Let $\mu_{g}:=\sqrt{\det g_{ij}}$ denote the Riemannian volume density. The evolution of $\mu_{g}$ is governed by:
$\partial_{T}\mu_{g}=\frac{1}{2}\mu_{g}\operatorname{tr}_{g}(\partial_{T}g)=-n\mu_{g}\left(1-\frac{N}{n}\right),$
which implies
$\mu_{g}(T)=\mu_{g}(T_{0})\exp\left(n\int_{T_{0}}^{T}\left(\frac{N}{n}-1\right)(s)\,ds\right).$
Applying the bootstrap assumption gives:
$\left|\log\left(\frac{\mu_{g}(T)}{\mu_{g}(T_{0})}\right)\right|\leq n\int_{T_{0}}^{T}\left|\frac{N}{n}-1\right|(s)ds\leq\frac{n\mathds{L}}{2\gamma}e^{-2\gamma T_{0}}.$
Exponentiating, we obtain the bounds:
$\mu_{g}(T_{0})e^{-\frac{n\mathds{L}}{2\gamma}e^{-2\gamma T_{0}}}\leq\mu_{g}(T)\leq\mu_{g}(T_{0})e^{\frac{n\mathds{L}}{2\gamma}e^{-2\gamma T_{0}}},$
or
$|\mu_{g}(T)-\mu_{g}(T_{0})|\leq\mu_{g}(T_{0})|e^{\frac{n\mathds{L}}{2\gamma}e^{-2\gamma T_{0}}}-1|\lesssim\mathds{L}e^{-2\gamma T_{0}}.$
which imply, in particular, the uniform bounds on $\det g_{ij}$. The constants $C_{1},C_{2}$ depend only on the initial data. ∎
Corollary 3.1 (Exponential Stability of Volume Function).
Under the same assumptions as in Proposition 3.1, the total volume functional satisfies the bound:
$\left|\operatorname{Vol}(g(T))-\operatorname{Vol}(g(T_{0}))\right|\lesssim\mathds{L}e^{-2\gamma T_{0}},\qquad\text{for all }T\in[T_{0},T_{\infty}],$
where $\operatorname{Vol}(g(T))=\int_{M}\mu_{g}(T)\,dx$ denotes the total volume with respect to $g(T)$.
Proof.
From the previous proposition and the evolution equation for $\mu_{g}$, we compute:
$\frac{d}{dT}\operatorname{Vol}(g(T))=\int_{M}\partial_{T}\mu_{g}\,dx=n\int_{M}\mu_{g}\left(\frac{N}{n}-1\right)dx.$
Integrating in time and using the uniform bound $\left\|\frac{N}{n}-1\right\|_{L^{\infty}(M)}\leq\mathds{L}e^{-2\gamma T}$, we deduce:
$\left|\operatorname{Vol}(g(T))-\operatorname{Vol}(g(T_{0}))\right|\lesssim\operatorname{Vol}(g(T_{0}))\int_{T_{0}}^{T}\mathds{L}e^{-2\gamma t}dt\lesssim\mathds{L}e^{-2\gamma T_{0}}.$
This completes the proof. ∎
Proposition 3.2.
Let $(x,T)\in M\times[T_{0},T_{\infty}]$ denote coordinates on a globally hyperbolic spacetime foliated by constant mean curvature (CMC) hypersurfaces, and let $g(T,x)$ denote the induced Riemannian metric on the CMC slices in CMC-transported coordinates. Denote by $\alpha(T)$ and $\beta(T)$ the maximal and minimal eigenvalues, respectively, of the symmetric bilinear form $g(T,x)$ with respect to the initial metric $g(T_{0},x)$. Then under the bootstrap assumption (87), we have the estimate
$|\alpha(T)-1|+|\beta(T)-1|\lesssim\mathds{L}e^{-2\gamma T_{0}},$
where $\mathds{L}$ is a constant depending on the norms in the bootstrap assumptions and $\gamma>0$ is a fixed decay rate.
Proof.
Fix a point $x\in M$, and consider the symmetric bilinear form $g(T,x)$ on $T_{x}M$. Define the maximal and minimal eigenvalues of $g(T,x)$ with respect to $g(T_{0},x)$ at the point $x$ as
(94) $\alpha(x,T):=\sup_{0\neq Y\in T_{x}M}\frac{g(T,x)(Y,Y)}{g(T_{0},x)(Y,Y)},\quad\beta(x,T):=\inf_{0\neq Y\in T_{x}M}\frac{g(T,x)(Y,Y)}{g(T_{0},x)(Y,Y)}.$
We aim to estimate $|\alpha(x,T)-1|$ and $|\beta(x,T)-1|$ uniformly in $x$ and $T\in[T_{0},T_{\infty}]$. Recall the evolution equation for the metric in CMC-transported coordinates:
(95) $\partial_{T}g_{ij}=-\frac{2\varphi}{\tau}N\Sigma_{ij}-2\left(1-\frac{N}{n}\right)g_{ij},$
where $N$ is the lapse function, $\Sigma_{ij}$ is the transverse-traceless part of the second fundamental form, $\tau$ is the mean curvature, and $\varphi$ is a smooth weight function. For a fixed tangent vector $Y\in T_{x}M$, we obtain from (171)
(96) $\partial_{T}g(Y,Y)=-\frac{2\varphi}{\tau}N\Sigma(Y,Y)-2\left(1-\frac{N}{n}\right)g(Y,Y).$
To estimate the right-hand side of (96), we begin with the bound
(97) $|\Sigma(Y,Y)|\leq\sqrt{g(T_{0},x)^{ik}g(T_{0},x)^{jl}\Sigma_{ij}\Sigma_{kl}}\cdot g(T_{0},x)(Y,Y),$
since $\Sigma$ is a symmetric tensor and $g(T_{0},x)$ is a fixed reference metric. Thus, using the bootstrap assumption (87), we have
$|\Sigma(Y,Y)|\lesssim\|\Sigma\|_{L^{\infty}(M)}\cdot g(T_{0},x)(Y,Y).$
Now, using (96), we integrate from $T_{0}$ to $T$:
(98) $|g(T,x)(Y,Y)-g(T_{0},x)(Y,Y)|$ $\leq\int_{T_{0}}^{T}\left|\partial_{s}g(s,x)(Y,Y)\right|ds$
(99) $\lesssim\int_{T_{0}}^{T}\left(\left|\frac{\varphi(s)}{\tau(s)}\right|\cdot|N\Sigma(Y,Y)|+\left|1-\frac{N}{n}\right|g(Y,Y)\right)ds.$
From the bootstrap assumption, we have the decay bounds:
$\left|\frac{\varphi(s)}{\tau(s)}\right|\lesssim e^{-s},\quad\|\Sigma\|_{L^{\infty}(M)}\lesssim e^{-\gamma s},\quad\left\|1-\frac{N}{n}\right\|_{L^{\infty}(M)}\lesssim e^{-2\gamma s}.$
Therefore,
(100) $|g(T,x)(Y,Y)-g(T_{0},x)(Y,Y)|$ $\lesssim\int_{T_{0}}^{T}e^{-s}e^{-\gamma s}g(T_{0},x)(Y,Y)ds+\int_{T_{0}}^{T}e^{-\gamma s}g(T_{0},x)(Y,Y)ds$
(101) $\lesssim\left(\int_{T_{0}}^{\infty}e^{-(1+\gamma)s}ds+\int_{T_{0}}^{\infty}e^{-2\gamma s}ds\right)g(T_{0},x)(Y,Y).$
This yields
$|g(T,x)(Y,Y)-g(T_{0},x)(Y,Y)|\lesssim e^{-2\gamma T_{0}}g(T_{0},x)(Y,Y).$
To improve this to an estimate of relative deviation, we write
$\left|\frac{g(T,x)(Y,Y)}{g(T_{0},x)(Y,Y)}-1\right|\lesssim e^{-2\gamma T_{0}}.$
Taking the supremum and infimum over unit vectors $Y$ with respect to $g(T_{0},x)$ as in (94), we obtain
$|\alpha(x,T)-1|+|\beta(x,T)-1|\lesssim\mathds{L}e^{-2\gamma T_{0}}.$
Since the estimates are uniform in $x$, the result follows. ∎
Remark 8.
Propositions 3.1 and 3.2 establish uniform two-sided bounds on the evolving spatial metric $g(T,x)$ in terms of the initial metric $g_{0}:=g(T_{0},x)$, valid on the entire bootstrap interval $T\in[T_{0},T_{\infty}]$, with $T_{0}\gg 1$ sufficiently large. More precisely, there exists a numerical constant $C\geq 1$, such that
(102) $C^{-1}g_{0}\leq g(T,x)\leq Cg_{0},$
in the sense of positive-definite symmetric bilinear forms. In particular, the evolving metric remains uniformly equivalent to the initial metric along the foliation. This uniform equivalence enables the definition of time-dependent Sobolev norms $\|\cdot\|_{H^{s}(g(T))}$ for tensor fields on the spatial slices $M_{T}$, and ensures that all Sobolev inequalities, elliptic estimates, derived from the geometry of $g_{0}$ continue to hold with constants depending only on $C$. This equivalence plays a critical role in establishing quantitative a priori estimates.
We prove the estimates on the Sobolev constants in the following section.
3.5. Controlling the Sobolev Constants
One of the fundamental aspects of this problem is to control the appropriate Sobolev norms of the solution variables ($\Sigma,\mathfrak{T},N)$ for large times. In addition, one needs to make use of the Sobolev inequalities for the tensor fields on a manifold. Recall that the Sobolev norms are defined with respect to the dynamical metric $g$. This metric verifies the transport equation 64. Therefore, one needs to make sure that the metric $g$ does not degenerate during the evolution. Using the estimates derived on the metric components, we define the isoperimetric constant for the Cauchy slice
(103) $\text{I}(M):=\sup_{U\subset M,\partial U\in C^{1}}\frac{\min\{\text{vol}(U),\text{vol}(U^{c})\}}{[\text{Area}(\partial U)]^{\frac{3}{2}}}.$
The following proposition yields an upper bound for $I(M)$
Proposition 3.3.
Let $(\mathcal{M},g)$ be a globally hyperbolic spacetime endowed with a smooth time function $T\colon\mathcal{M}\to\mathbb{R}$ whose level sets
$M_{T}:=\{T=\mathrm{const}\}$
form a foliation by compact $2$-dimensional Riemannian manifolds (with induced metric still denoted $g(T)$). Assume the initial data on $M_{T_{0}}$ satisfy the hypotheses of Section 2 and that the bootstrap assumption (2.10) holds along the entire slab $T\in[T_{0},T_{\infty})$. Then for every $T\in[T_{0},T_{\infty})$ the isoperimetric constant satisfies
(104) $I(M_{T})\;\leq\;10\,I(M_{T_{0}}).$
Proof.
Fix any $T\in[T_{0},T_{\infty})$ and let $U\subset M_{T}$ be an arbitrary Caccioppoli set whose boundary $\partial U$ is of class $C^{1}$. Let $\Phi_{s}$ denote the flow generated by the future-directed vector field $\partial_{T}$, so that $\Phi_{T-T_{0}}\colon M_{T_{0}}\to M_{T}$ is a diffeomorphism. Set
$U_{0}:=\Phi_{-(T-T_{0})}(U)\subset M_{T_{0}},\qquad U^{c}_{0}:=\Phi_{-(T-T_{0})}(U^{c}).$
By global hyperbolicity, the integral curves of $\partial_{T}$ intersect each $M_{T}$ exactly once, so the map $\Phi_{T-T_{0}}$ is globally well-defined and smooth.
Let $\nu$ denote the outward unit normal to $\partial U$ in $(M_{T},g(T))$ and let $\nu_{0}$ denote the outward unit normal to $\partial U_{0}$ in $(M_{T_{0}},g(T_{0}))$. By Proposition 3.2, the differential $D\Phi_{T-T_{0}}|_{T_{x}(\partial U_{0})}$ satisfies the uniform ellipticity bounds
(105) $\beta(T)\;\leq\;\frac{d\sigma_{g(T)}}{d\sigma_{g(T_{0})}}\Big|_{\partial U}\;\leq\;\alpha(T),$
where $\beta(T)>0$ and $\alpha(T)<\infty$ are the eigenvalues as in Proposition 3.2 and verify the uniform estimates in the same proposition 3.2. In particular,
(106) $\mathrm{Area}_{g(T)}(\partial U)\;\geq\;\beta(T)\,\mathrm{Area}_{g(T_{0})}(\partial U_{0}).$
Similarly, Proposition 3.2 implies pointwise bounds for the Riemannian volume form:
(107) $\beta(T)^{\frac{3}{2}}\;\leq\;\frac{d\mathrm{vol}_{g(T)}}{d\mathrm{vol}_{g(T_{0})}}\;\leq\;\alpha(T)^{\frac{3}{2}},$
where $\alpha(T),\beta(T)$ are likewise controlled by 3.2. Integrating (107) over $U_{0}$ and $U^{c}_{0}$ yields
(108) $\mathrm{Vol}_{g(T)}(U)\;\leq\;\alpha(T)^{\frac{3}{2}}\,\mathrm{Vol}_{g(T_{0})}(U_{0}),\qquad\mathrm{Vol}_{g(T)}(U^{c})\;\leq\;\alpha(T)^{\frac{3}{2}}\,\mathrm{Vol}_{g(T_{0})}(U^{c}_{0}).$
By definition of the isoperimetric constant on $M_{T}$,
$I(M_{T})=\sup_{U\subset M,\partial U\in C^{1}}\frac{\min\{\text{vol}(U),\text{vol}(U^{c})\}}{[\text{Area}(\partial U)]^{\frac{3}{2}}}.$
Using (106) and (108) gives
$I(M_{T})\leq\frac{\alpha(T)^{\frac{3}{2}}}{\beta(T)^{\frac{3}{2}}}\sup_{U\subset M,\partial U_{0}\in C^{1}}\frac{\min\{\text{vol}(U_{0}),\text{vol}(U^{c}_{0})\}}{[\text{Area}(\partial U_{0})]^{\frac{3}{2}}}=\frac{\alpha(T)^{\frac{3}{2}}}{\beta(T)^{\frac{3}{2}}}I(M_{T_{0}}).$
Now in light of the proposition 3.2, the ratio $\frac{\alpha(T)^{\frac{3}{2}}}{\beta(T)^{\frac{3}{2}}}$ remains contained in the interval [1/10, 10][1/10,\,10] for all $T\in[T_{0},T_{\infty})$. Hence, $I(M_{T})\;\leq\;10\,I(M_{T_{0}})$, which establishes the desired estimate (104). ∎
Proposition 3.4 (Sobolev Embedding for Tensor Fields).
Let $(M,g)$ be a closed (i.e., compact without boundary), $C^{\infty}$ 33-dimensional Riemannian manifold, and let $\psi\in C^{\infty}(\Gamma({}^{K}TM\otimes{}^{L}T^{*}M))$ be a smooth section of a mixed tensor bundle over $M$. Then the following Sobolev-type inequalities hold:
[…]
where the constants are purely numerical.
Proof.
The proof proceeds via localization, partition of unity, and a scalar reduction argument adapted to tensor fields. First, since $M$ is a closed Riemannian manifold, it admits a finite atlas $\{(U_{i},\varphi_{i})\}_{i=1}^{N}$ such that each chart $\varphi_{i}:U_{i}\to\mathbb{R}^{3}$ is a diffeomorphism onto its image, and the pulled-back metric components $g_{ij}$ and their derivatives are uniformly bounded on each chart. Let $\{\chi_{i}\}_{i=1}^{N}$ be a smooth partition of unity subordinate to this cover.
Let $\psi\in C^{\infty}(\Gamma({}^{K}TM\otimes{}^{L}T^{*}M))$. Define for $\epsilon>0$ the scalar function:
$f_{\epsilon}:=\sqrt{|\psi|^{2}+\epsilon},$
where $|\psi|^{2}=g^{i_{1}j_{1}}\cdots g^{i_{K}j_{K}}g_{a_{1}b_{1}}\cdots g_{a_{L}b_{L}}\psi_{i_{1}\cdots i_{K}}^{a_{1}\cdots a_{L}}\psi_{j_{1}\cdots j_{K}}^{b_{1}\cdots b_{L}}$ is the pointwise norm squared induced by $g$.
Since the Levi-Civita connection $\nabla$ is compatible with $g$, we compute:
$\nabla f_{\epsilon}=\frac{1}{\sqrt{|\psi|^{2}+\epsilon}}\langle\psi,\nabla\psi\rangle_{g},$
where the inner product $\langle\psi,\nabla\psi\rangle_{g}$ is taken pointwise with respect to $g$. This yields the pointwise bound:
$|\nabla f_{\epsilon}|\leq\frac{|\psi|\cdot|\nabla\psi|}{\sqrt{|\psi|^{2}+\epsilon}}\leq|\nabla\psi|.$
We now apply the standard scale-invariant Sobolev inequality for scalar functions on 3-dimensional Riemannian manifolds:
[…]
Using the inequality $f_{\epsilon}||_{L^{p}(M)}\leq\||\psi|\|_{L^{p}(M)}+\epsilon^{1/2}|M|^{1/p}$ and passing to the limit as $\epsilon\to 0$ yields
[…]
This proves inequality (109).
We assume the scalar version of the scale-invariant Sobolev inequality:
[…]
for all smooth real-valued functions $f$. Let $\psi\in C^{\infty}(\Gamma({}^{K}TM\otimes{}^{L}T^{*}M))$ be a smooth mixed tensor field. Set
$f:=|\psi|_{g},$
the pointwise norm induced by $g$. Then $f$ is a smooth nonnegative scalar function, and
$\|\psi\|_{L^{\infty}}=\|f\|_{L^{\infty}},\qquad\|\psi\|_{L^{4}}=\|f\|_{L^{4}}.$
For tensor fields on a Riemannian manifold, one has the pointwise Kato inequality as seen in the proof of the first inequality (110)
(112) $|\nabla f|\leq|\nabla\psi|.$
This follows from writing $f=(\psi_{\alpha}\psi_{\alpha}+\epsilon)^{1/2}$ in a local chart and and applying the Cauchy-Schwarz inequality. Thus
$\|\nabla f\|_{L^{4}}\leq\|\nabla\psi\|_{L^{4}}.$
Applying (111) to the scalar function $f=|\psi|$ gives
$\|\psi\|_{L^{\infty}}=\|f\|_{L^{\infty}}\leq C\Big(\max(1,I(M,g))\Big)^{\frac{1}{2}}\Big({\rm Vol}_{g}(M)^{\frac{1}{3}}\|\nabla f\|_{L^{4}}+{\rm Vol}_{g}(M)^{-\frac{1}{4}}\|f\|_{L^{4}}\Big).$
Using Kato’s inequality (112) and the equalities $\|f\|_{L^{4}}=\|\psi\|_{L^{4}}$, $\|\nabla f\|_{L^{4}}\leq\|\nabla\psi\|_{L^{4}}$, we obtain
$\|\psi\|_{L^{\infty}}\leq C\Big(\max(1,I(M,g))\Big)^{\frac{1}{2}}\Big({\rm Vol}_{g}(M)^{\frac{1}{3}}\|\nabla\psi\|_{L^{4}}+{\rm Vol}_{g}(M)^{-\frac{1}{4}}\|\psi\|_{L^{4}}\Big),$
which is exactly the desired tensorial inequality (110). ∎
Corollary 3.2.
For the dynamical manifold $(M_{T},g(T))$, the Sobolev inequalities 109-110 hold with the constant $C\bigg(\max(1,I(M,g))\bigg)$ being uniformly bounded by $O(1)$ purely numerical entity $\forall T\in[T_{0},T_{\infty})$ in light of the proposition $3.$
Lastly, we note that on $(M,g)$ we can obtain compactness theorems for Sobolev embeddings. This can be proven by working on local charts using the uniform estimates on the metric $g(t)$ on the interval $T_{0},T_{\infty}]$ in light of propositions 3.1-3.2 and gluing the charts together in a compatible manner. In particular, on $(M,g)$ for a $C^{\infty}$ tensor field $\psi$, the Sobolev norm $H^{I}(M)$ is defined as
[…]
Proposition 3.5.
The embedding $H^{I}(M)\hookrightarrow H^{J}(M)$ is compact for $I>J$.
Theorem 3.1 (Local Well-Posedness and Continuation).
Let $g_{0}$ be a Riemannian metric on a compact 3-manifold $M$, satisfying the uniform ellipticity bounds
$C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0}$
for some smooth background metric $\xi_{0}$ and constant $C>0$. Suppose the initial data $(\Sigma_{0},\mathfrak{T}_{0})$ belong to the Sobolev space $H^{I}\times H^{I-1}$ for some $I>\frac{5}{2}$, and satisfy the constraint equations (67)-(68) associated with the re-scaled evolution system (74)–(75), supplemented by the elliptic lapse equation (76) in constant mean curvature (CMC) transported spatial gauge.
Then there exists a time $t^{*}>0$, depending only on $C$, $\|\Sigma_{0}\|_{H^{I}}$, and $\|\mathfrak{T}_{0}\|_{H^{I-1}}$, such that the Cauchy problem admits a unique solution
$(\Sigma(t),\mathfrak{T}(t))\in C\left([0,t^{*}];H^{I}\times H^{I-1}\right),$
with lapse $N(t)\in H^{I+2}$, and the solution map
$(\Sigma_{0},\mathfrak{T}_{0})\mapsto(\Sigma(t),\mathfrak{T}(t),N(t))$
is continuous as a map
$H^{I}\times H^{I-1}\to C\left([0,t^{*}];H^{I}\times H^{I-1}\times H^{I+2}\right).$
Let $t^{\max}\geq t^{*}$ denote the maximal time of existence of the solution. Then either $t^{\max}=\infty$, or the solution breaks down at $t^{\max}$ in the sense that
[…]
where $\{\mu_{i}(t)\}_{i=1}^{3}$ denote the eigenvalues of the symmetric endomorphism $\xi_{0}^{-1}g(t)$.
4. Estimates
In this section, we derive uniform estimates for the spatial Riemann curvature tensor $\operatorname{Riem}_{ijkl}$ on each constant mean curvature (CMC) slice, under the bootstrap assumptions introduced in Section 3.3. Our goal is to express $\operatorname{Riem}_{ijkl}$ entirely in terms of the renormalized dynamical variables, namely the trace-free part of the second fundamental form $\Sigma$ and the auxiliary tensor $\mathfrak{T}$, and to obtain estimates compatible with the Sobolev regularity used in our energy framework.
To this end, we recall that in three spatial dimensions, the Riemann curvature tensor admits the following Ricci decomposition:
(114) $\operatorname{Riem}_{ijkl}=\frac{1}{2}\left(g_{ik}\operatorname{Ric}_{jl}+g_{jl}\operatorname{Ric}_{ik}-g_{il}\operatorname{Ric}_{jk}-g_{jk}\operatorname{Ric}_{il}\right)+\frac{1}{6}R(g)\left(g_{il}g_{jk}-g_{ik}g_{jl}\right),$
where $\operatorname{Ric}_{ij}$ and $R(g)$ denote, respectively, the Ricci tensor and scalar curvature of the induced Riemannian metric $g$ on the CMC slice.
The Ricci tensor can be expressed algebraically in terms of the renormalized momentum variable $\mathfrak{T}$ via the relation
(115) $\operatorname{Ric}_{ij}=\mathfrak{T}_{ij}+\frac{1}{n}R(g)g_{ij},$
Furthermore, the scalar curvature satisfies the renormalized Hamiltonian constraint, which reads
(116) $R(g)+\frac{n-1}{n}=|\Sigma|_{g}^{2},$
where $|\Sigma|_{g}^{2}=g^{ia}g^{jb}\Sigma_{ij}\Sigma_{ab}$ denotes the pointwise squared norm of the symmetric trace-free tensor $\Sigma$.
Substituting (115) and (116) into (114), we conclude that the full spatial Riemann curvature tensor $\operatorname{Riem}_{ijkl}$ is entirely determined by the tensors $\mathfrak{T}$ and $\Sigma$, together with the metric $g$. That is,
(117) $\operatorname{Riem}_{ijkl}=\mathcal{F}_{ijkl}(g,\Sigma,\mathfrak{T}),$
where $\mathcal{F}_{ijkl}$ denotes a universal algebraic expression involving the metric, its inverse, and the tensors $\mathfrak{T}$ and $\Sigma$. In particular, any Sobolev or pointwise bound on $\mathfrak{T}$ and $\Sigma$ yields a corresponding bound on the full curvature tensor.
This observation is central to the curvature estimates employed throughout the remainder of the analysis. In particular, it ensures that the control of geometric quantities such as $\operatorname{Riem}$ reduces to the control of the dynamical variables governed by the Einstein evolution equations in the chosen CMC-transported spatial gauge.  [^3] In $n\geq 4$, one needs to estimate the additional Weyl curvature as well. In such case, one first imposes bootstrap assumption on Weyl curvature and then improves the bootstrap by means of separate energy estimates for the Weyl curvature.
4.1. Elliptic estimates
Recall the elliptic equation
(118) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
We obtain the following estimate for the normalized lapse function $\frac{N}{n}-1$
Proposition 4.1.
Let $(M,g)$ be a compact Riemannian $3$-manifold with metric $g$ satisfying the bootstrap assumption (87), and let $\Sigma$ denote a symmetric transverse-traceless (TT) $2$-tensor on $M$. Consider the lapse function $N$ solving the elliptic equation
$\Delta_{g}\left(\frac{N}{n}-1\right)+\left(|\Sigma|_{g}^{2}+\frac{1}{n}\right)\left(\frac{N}{n}-1\right)=-|\Sigma|_{g}^{2},$
where $n=3$. Then for any integer $I\in\{2,3\}$, the following Sobolev estimate holds:
(119) $\sum_{j=0}^{I+2}\|\nabla^{j}\big(\frac{N}{n}-1\big)\|_{L^{2}(M)}\lesssim(1+\Gamma)\sum_{j=0}^{I}\|\nabla^{j}\Sigma\|_{L^{2}(M)}^{2},$
where $\nabla$ denotes the Levi-Civita connection of $g$ and $\Gamma$ controls curvature norms as specified in the bootstrap regime.
Remark 9.
Note the loss of decay of $(\frac{N}{n}-1)$ at the top-most order due to loss of decay of $\Sigma$ at the top-most order.
Proof.
Under the bootstrap assumption (87), we establish energy estimates for the spatial Riemann curvature tensor $\operatorname{Riem}$ of the Riemannian metric $g$ on the spatial slice $M$. More precisely, for all integers $0\leq I\leq 2$, we claim the following uniform bound:
(120) $\|\nabla^{I}\operatorname{Riem}\|_{L^{2}(M)}\lesssim\Gamma+1.$
To derive this, we exploit the classical Ricci decomposition of the Riemann tensor in dimension $n=3$, where the Weyl tensor vanishes identically:
(121) $\operatorname{Riem}_{ijkl}=\tfrac{1}{2}\left(g_{ik}\operatorname{Ric}_{jl}+g_{il}\operatorname{Ric}_{kj}-g_{jk}\operatorname{Ric}_{il}-g_{jl}\operatorname{Ric}_{ik}\right)+\tfrac{1}{6}R(g)\left(g_{il}g_{jk}-g_{ik}g_{jl}\right).$
From this decomposition, it follows that the full Riemann tensor is algebraically expressible in terms of the Ricci tensor and scalar curvature. Invoking the renormalized Hamiltonian constraint, we write:
(122) $R(g)=-\tfrac{n-1}{n}+|\Sigma|^{2},$
where $\Sigma$ denotes the trace-free part of the second fundamental form in CMC gauge, and where all norms and derivatives are computed with respect to the spatial metric $g$.
Let us now provide schematic bounds for $\nabla^{I}\operatorname{Riem}$ in terms of the fundamental dynamical fields $\mathfrak{T}$ and $\Sigma$. Using the expressions for $\operatorname{Ric}$ and $R(g)$, we obtain:
(123) $|\nabla\operatorname{Riem}|$ $\lesssim|\nabla\mathfrak{T}|+|\nabla R(g)|\lesssim|\nabla\mathfrak{T}|+|\Sigma||\nabla\Sigma|,$
(124) $|\nabla^{2}\operatorname{Riem}|$ $\lesssim|\nabla^{2}\mathfrak{T}|+|\nabla\Sigma|^{2}+|\Sigma||\nabla^{2}\Sigma|.$
By applying the Cauchy–Schwarz inequality and standard Sobolev embedding on the compact manifold $M$, together with the bootstrap assumption and bounds derived in Propositions 3.13.2, we deduce:
(125) $\|\nabla\operatorname{Riem}\|_{L^{2}(M)}$ $\lesssim\|\nabla\mathfrak{T}\|_{L^{2}(M)}+\|\nabla\Sigma\|_{L^{2}(M)}\|\nabla^{2}\Sigma\|_{L^{2}(M)}\lesssim\Gamma+e^{-2\gamma T}\mathds{Y}^{2}\lesssim\Gamma+1,$
(126) $\|\nabla^{2}\operatorname{Riem}\|_{L^{2}(M)}$ $\lesssim\|\nabla^{2}\mathfrak{T}\|_{L^{2}(M)}+\|\nabla^{2}\Sigma\|_{L^{2}(M)}^{2}\lesssim\Gamma+e^{-2\gamma T}\mathds{Y}^{2}\lesssim\Gamma+1.$
Finally, the $L^{2}$ bound for $\operatorname{Riem}$ without derivatives follows directly from the pointwise Ricci decomposition and the bootstrap bound on $\mathfrak{T}$ and $\Sigma$:
(127) $\|\operatorname{Riem}\|_{L^{2}(M)}\lesssim\Gamma+1.$
Thus, we conclude that all desired curvature norms up to three derivatives are uniformly controlled by the bootstrap parameter $\Gamma$, thereby validating the estimate (120). Now we are ready to obtain the estimates for $\frac{N}{n}-1$. Let us recall the elliptic equation
(128) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
Let us denote the operator $\Delta_{g}+(|\Sigma|^{2}+\frac{1}{n})$ by $\mathscr{L}$. We claim that an estimate of the following type holds for $u\in C^{\infty}(M)$
[…]
with the regularity $\nabla^{I-1}\text{Riem}\in L^{2}(M)$. To do this we commute the lapse equation with $\nabla^{I}$ for $0\leq I\leq 3$
$\mathscr{L}\nabla^{I}(\frac{N}{n}-1)=-\Delta_{g}\nabla^{I}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})\nabla^{I}(\frac{N}{n}-1)=\nabla^{I}\mathscr{L}(\frac{N}{n}-1)$
$+\sum_{J_{1}+J_{2}+J_{3}=I-1}\nabla^{J_{1}+1}\Sigma\nabla^{J_{2}}\Sigma\nabla^{J_{3}}(\frac{N}{n}-1)+\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\nabla^{J_{1}}\text{Ric}\nabla^{J_{2}+I-m}(\frac{N}{n}-1)$
$+\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\nabla^{J_{1}}\text{Riem}\nabla^{J_{2}+I-m}(\frac{N}{n}-1)+\sum_{m=0}^{I-2}\sum_{J_{1}+J_{2}=m}\nabla^{J_{1}+1}\text{Riem}\nabla^{J_{2}+I-m-1}(\frac{N}{n}-1).$
The elliptic regularity implies
[…]
Under the boot-strap assumption 87, we estimate each term separately. First estimate
[…]
The next term reads
[…]
for the following terms we have
[…]
Therefore collecting all the terms together, we conclude
[…]
Now we prove the desired estimates in an iterative way starting from the first-order estimate. First, recall the elliptic equation
(133) $-\Delta_{g}(\frac{N}{n}-1)+(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)=-|\Sigma|^{2}.$
Now multiply both sides by $\frac{N}{n}-1$ and integrate by parts
(134) $\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}(|\Sigma|^{2}+\frac{1}{n})(\frac{N}{n}-1)^{2}\mu_{g}=-\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)\mu_{g}.$
Now we estimate the term on the right-hand side by Cauchy-Schwartz as follows
(135) $-\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)\mu_{g}\leq(n+1)\int_{M}|\Sigma|^{4}\mu_{g}+\frac{1}{n+1}\int_{M}(\frac{N}{n}-1)^{2}\mu_{g}.$
Therefore, 134 reads
$\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)^{2}\mu_{g}+(\frac{1}{n}-\frac{1}{n+1})\int_{M}(\frac{N}{n}-1)^{2}\mu_{g}\leq(n+1)\int_{M}|\Sigma|^{4}\mu_{g}$
which by means of the inequality 3.4 yields
$\int_{M}|\nabla(\frac{N}{n}-1)|^{2}\mu_{g}+\int_{M}|\Sigma|^{2}(\frac{N}{n}-1)^{2}\mu_{g}+\frac{1}{n(n+1)}\int_{M}(\frac{N}{n}-1)^{2}\mu_{g}$
$\lesssim\left(\int_{M}|\nabla\Sigma|^{2}\mu_{g}+\int_{M}|\Sigma|^{2}\mu_{g}\right)^{2}.$
More concretely, we have the two lowest-order estimates
[…]
Now iterate this using 132 to obtain
$||\frac{N}{n}-1||_{H^{2}(M)}\lesssim|||\Sigma|^{2}||_{L^{2}(M)}+(1+\Gamma)||\frac{N}{n}-1)||_{L^{2}(M)}\lesssim|||\Sigma|^{2}||_{L^{2}(M)}+(1+\Gamma)||\Sigma||^{2}_{H^{1}(M)},$
$||\frac{N}{n}-1||_{H^{3}(M)}\lesssim|||\Sigma|^{2}||_{H^{1}(M)}+(1+\Gamma)||\frac{N}{n}-1)||_{H^{1}(M)}\lesssim|||\Sigma|^{2}||_{H^{1}(M)}+(1+\Gamma)||\Sigma||^{2}_{H^{1}(M)},$
$||\frac{N}{n}-1||_{H^{4}(M)}\lesssim|||\Sigma|^{2}||_{H^{2}(M)}+(1+\Gamma)||\frac{N}{n}-1)||_{H^{2}(M)}\lesssim|||\Sigma|^{2}||_{H^{2}(M)}+(1+\Gamma)||\Sigma||^{2}_{H^{1}(M)}$
and so on. ow use the embedding $H^{i_{1}}(M)\hookrightarrow H^{i_{2}}(M)$ for $i_{1}>i_{2}$, and the algebra property of the Sobolev spaces $H^{s}(M)$ for $s>\frac{n}{2}$ or that of $L^{\infty}(M)\cap H^{s}(M)$ for $s\geq 0$, we conclude the desired estimate
[…]
The elliptic estimate yields, together with the algebra property of Sobolev spaces for $I>2$
[…]
The next part entails controlling the lower order norm $\mathcal{F}$ of $\Sigma$ in terms of the top order norm $\mathcal{O}$ and the initial data norm $\mathcal{I}^{0}$.
4.2. Estimating the lower order norm of TT tensor $\Sigma$ via transport equation for $\Sigma$
In this section, we prove that the weighted lower order norm of the $TT$ tensor $\Sigma$ is controlled by the top order norm of $(\mathfrak{T},\Sigma)$ and the initial data. Recall the equation for $\Sigma$
(140) $\frac{\partial\Sigma_{ij}}{\partial T}$ $=$ $-(n-1)\Sigma_{ij}-\frac{\varphi}{\tau}N\mathfrak{T}_{ij}+\frac{\varphi}{\tau}\nabla_{i}\nabla_{j}(\frac{N}{n}-1)+\frac{2\varphi}{\tau}N\Sigma_{ik}\Sigma^{k}_{j}$
$-\frac{\varphi}{n\tau}(\frac{N}{n}-1)g_{ij}-(n-2)(\frac{N}{n}-1)\Sigma_{ij}.$
To estimate the lower order norm $\mathcal{F}$, we will treat the equation for $\Sigma$ as a transport equation.
Definition 6.
(141) $\mathscr{E}^{low}:=\frac{1}{2}\sum_{I\leq 2}\int_{M}|\nabla^{I}\Sigma|^{2}\mu_{g}$
First note that $\mathcal{F}^{2}\approx e^{2\gamma T}\mathscr{E}^{low}$. We have the following theorem estimating $\mathcal{F}^{2}$.
Theorem 4.1 (Coercive control of the reversed entropy).
Let $(M^{n},g(T))$ be a smooth solution of the rescaled Einstein vacuum equations with cosmological constant $\Lambda>0$ on a time interval $T_{0},T]$, written in the fixed gauge of this work. Let $\mathscr{E}^{\mathrm{low}}(T)\geq 0$ denote the reversed entropy functional controlling the lower order norm of $\Sigma$ as defined in 6 and let $\mathcal{F}(T)$ be the associated re-scaled entity i.e.,
(142) $e^{2T}\,\mathscr{E}^{\mathrm{low}}(T)\ \lesssim\ \mathcal{F}(T)^{2}\ \lesssim\ e^{2T}\,\mathscr{E}^{\mathrm{low}}(T)$
and $\mathcal{O}$ be the top order norm 84 Then the following estimate for the re-scaled reversed entropy holds
  1. (i)
    (A–priori upper bound).
    (143) $\mathcal{F}(T)^{2}\ \lesssim\ \Bigl(1+\bigl(\mathcal{I}^{0}\bigr)^{2}+\mathcal{O}^{2}\Bigr).$
In particular, combining (142) and (143) yields
(144) $e^{2T}\,\mathscr{E}^{\mathrm{low}}(T)\ \lesssim\ \Bigl(1+\bigl(\mathcal{I}^{0}\bigr)^{2}+\mathcal{O}^{2}\Bigr),$
where the involved constants are of pure numerical nature i.e., dpendend on the dimension and independent of $T$.
Proof.
Recall definition 6
(145) $\mathscr{E}^{low}(T):=\frac{1}{2}\sum_{I\leq 2}\int_{M(T)}\langle\nabla^{I}\Sigma,\nabla^{I}\Sigma\rangle\mu_{g}.$
Explicit application of the time evolution operator $\frac{d}{dT}$ together with the transport theorem [32] yields
(146) $\frac{d\mathscr{E}^{low}}{dT}=\sum_{I\leq 2}\int_{M}\langle\nabla^{I}\partial_{T}\Sigma,\nabla^{I}\Sigma\rangle\mu_{g}+\sum_{I\leq 2}\int_{M}\langle[\partial_{T},\nabla^{I}]\Sigma,\nabla^{I}\Sigma\rangle\mu_{g}+\mathscr{ER}_{2},$
where the error terms $\mathscr{ER}_{2}$ reads schematically
(147) $\mathscr{ER}_{2}\sim\sum_{I\leq 2}\int_{M(T)}\partial_{T}(g^{-1}....g^{-1}g.....g)\nabla^{I}\Sigma\nabla^{I}\Sigma\mu_{g}+\sum_{I\leq 2}\int_{M(T)}|\nabla^{I}\Sigma|^{2}\partial_{T}\mu_{g}.$
Using the equation for $\partial_{T}\Sigma$, $\frac{d\mathscr{E}^{low}}{dT}$ reads
$\frac{d\mathscr{E}^{low}}{dT}=-2(n-1)\mathscr{E}+\sum_{I\leq 2}\frac{\varphi}{\tau}\int_{M(T)}N\nabla^{I}\mathfrak{T}\nabla^{I}\Sigma\mu_{g}+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{M(T)}\frac{\varphi}{\tau}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\Sigma\nabla^{I}\Sigma\mu_{g}$
$+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}+J_{3}=m}\int_{M(T)}\frac{\varphi}{\tau}\nabla^{J_{1}}N\nabla^{J_{2}+1}\Sigma\nabla^{J_{3}+I-m-1}\Sigma\nabla^{I}\Sigma\mu_{g}+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\Sigma\nabla^{I}\Sigma\mu_{g}$
$+\frac{\varphi}{\tau}\sum_{I\leq 2}\sum_{J_{1}+J_{2}+J_{3}=I}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}}\Sigma\nabla^{J_{3}}\Sigma\nabla^{I}\Sigma\mu_{g}+\sum_{I\leq 2}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{I+2}(\frac{N}{n}-1)\nabla^{I}\Sigma\mu_{g}$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{M(T)}\nabla^{J_{1}}(\frac{N}{n}-1)\nabla^{J_{2}}\Sigma\nabla^{I}\Sigma\mu_{g}+\sum_{I\leq 2}\frac{\varphi}{\tau}\int_{M}\sum_{J_{1}+J_{2}=I-1}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}}\mathfrak{T}\nabla^{I}\Sigma\mu_{g}+\mathscr{ER}_{2}.$
[…]
(148) $|\sum_{I\leq 2}\frac{\varphi}{\tau}\int_{M(T)}N\nabla^{I}\mathfrak{T}\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-T}\mathcal{O}^{2}.$
The next term is estimated as follows
(149) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{M(T)}\frac{\varphi}{\tau}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\Sigma\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-(1+4\gamma)T}\mathds{Y}^{4},$
(150) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}+J_{3}=m}\int_{M(T)}\frac{\varphi}{\tau}\nabla^{J_{1}}N\nabla^{J_{2}+1}\Sigma\nabla^{J_{3}+I-m-1}\Sigma\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-(1+3\gamma)T}\mathds{Y}^{3},$
(151) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\Sigma\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-4\gamma T}\mathds{Y}^{4},$
(152) $|\frac{\varphi}{\tau}\sum_{I\leq 2}\sum_{J_{1}+J_{2}+J_{3}=I}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}}\Sigma\nabla^{J_{3}}\Sigma\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-(1+3\gamma)T}\mathds{Y}^{3},$
(153) $|\sum_{I\leq 2}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{I+2}(\frac{N}{n}-1)\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-(1+3\gamma)T}\mathds{Y}^{3},$
(154) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{M(T)}\nabla^{J_{1}}(\frac{N}{n}-1)\nabla^{J_{2}}\Sigma\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-4\gamma T}\mathds{Y}^{4},$
(155) $|\sum_{I\leq 2}\frac{\varphi}{\tau}\int_{M}\sum_{J_{1}+J_{2}=I-1}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}}\mathfrak{T}\nabla^{I}\Sigma\mu_{g}|\lesssim e^{-(1+3\gamma)T}\Gamma\mathds{Y}^{3}.$
Collection of all terms yields
(156) $\frac{d\mathscr{E}}{dT}+2(n-1)\mathscr{E}\lesssim e^{-T}\mathcal{O}^{2}+e^{-(1+4\gamma)T}\mathds{Y}^{4}+e^{-(1+3\gamma)T}\mathds{Y}^{3}$
$+e^{-4\gamma T}\mathds{Y}^{4}+e^{-(1+3\gamma)T}\Gamma\mathds{Y}^{3}$
integration of which yields
$\mathscr{E}(T)\lesssim\;$ $e^{-2(n-1)(T-T_{0})}\mathscr{E}(T_{0)}$
$+e^{-2(n-1)T}\left(e^{[2(n-1)-1]T}-e^{[2(n-1)-1]T_{0}}\right)\mathcal{O}^{2}$
$+e^{-2(n-1)T}\left(e^{[2(n-1)-(1+4\gamma)]T}-e^{[2(n-1)-(1+4\gamma)]T_{0}}\right)\mathds{Y}^{4}$
$+e^{-2(n-1)T}\left(e^{[2(n-1)-(1+3\gamma)]T}-e^{[2(n-1)-(1+3\gamma)]T_{0}}\right)\mathds{Y}^{3}$
$+e^{-2(n-1)T}\left(e^{[2(n-1)-4\gamma]T}-e^{[2(n-1)-4\gamma]T_{0}}\right)\mathds{Y}^{4}$
(157) $+e^{-2(n-1)T}\left(e^{[2(n-1)-(1+3\gamma)]T}-e^{[2(n-1)-(1+3\gamma)]T_{0}}\right)\Gamma\mathds{Y}^{3}.$
Alternatively, we can write:
$e^{2\gamma T}\mathscr{E}(T)\leq\;$ $\frac{e^{2(n-1-\gamma)T_{0}}}{e^{2(n-1-\gamma)T}}e^{2\gamma T_{0}}\mathscr{E}(T_{0})$
$+e^{-2(n-1-\gamma)T}\left(e^{[2(n-1)-1]T}-e^{[2(n-1)-1]T_{0}}\right)\mathcal{O}^{2}$
$+e^{-2(n-1-\gamma)T}\left(e^{[2(n-1)-(1+4\gamma)]T}-e^{[2(n-1)-(1+4\gamma)]T_{0}}\right)\mathds{Y}^{4}$
$+e^{-2(n-1-\gamma)T}\left(e^{[2(n-1)-(1+3\gamma)]T}-e^{[2(n-1)-(1+3\gamma)]T_{0}}\right)\mathds{Y}^{3}$
$+e^{-2(n-1-\gamma)T}\left(e^{[2(n-1)-4\gamma]T}-e^{[2(n-1)-4\gamma)]T_{0}}\right)\mathds{Y}^{4}$
$=\;$ $\frac{e^{2(n-1-\gamma)T_{0}}}{e^{2(n-1-\gamma)T}}e^{2\gamma T_{0}}\mathscr{E}(T_{0})$
$+e^{-(1-2\gamma)T}\left(1-\frac{e^{[2(n-1)-1]T_{0}}}{e^{[2(n-1)-1]T}}\right)\mathcal{O}^{2}$
$+e^{-(1+2\gamma)T}\left(1-\frac{e^{[2(n-1)-(1+4\gamma)]T_{0}}}{e^{[2(n-1)-(1+4\gamma)]T}}\right)\mathds{Y}^{4}$
$+e^{-(1+\gamma)T}\left(1-\frac{e^{[2(n-1)-(1+3\gamma)]T_{0}}}{e^{[2(n-1)-(1+3\gamma)]T}}\right)\mathds{Y}^{3}$
$+e^{-2\gamma T}\left(1-\frac{e^{[2(n-1)-4\gamma]T_{0}}}{e^{[2(n-1)-4\gamma]T}}\right)\mathds{Y}^{3}$
$\lesssim\;$ $e^{2\gamma T_{0}}\mathscr{E}(T_{0})+e^{-(1-2\gamma)T}\mathcal{O}^{2}+e^{-(1+2\gamma)T_{0}}\mathds{Y}^{4}+e^{-(1+\gamma)T_{0}}\mathds{Y}^{3}+e^{-2\gamma T_{0}}\mathds{Y}^{3}$
(158) $\lesssim\;$ $e^{-(1-2\gamma)T}\mathcal{O}^{2}+(\mathcal{I}^{0})^{2}+1.$
We begin from this suboptimal decay estimate
(159) $e^{2\gamma T}\mathscr{E}^{low}(T):=\frac{e^{2\gamma T}}{2}\sum_{I\leq 2}\int_{M(T)}\langle\nabla^{I}\Sigma,\nabla^{I}\Sigma\rangle\,\mu_{g}\lesssim e^{-(1-2\gamma)T}\mathcal{O}^{2}+(\mathcal{I}^{0})^{2}+1,\forall T\in[T_{0},T_{\infty}].$
Observe that the bootstrap assumption (87) allows one to control $e^{2\gamma T}\mathscr{E}^{low}(T)$ uniformly by a constant multiple of $(\mathcal{I}^{0})^{2}+1$ provided that $\gamma\leq\frac{1}{2}$. We refrain from immediately fixing $\gamma$ at this threshold. Instead, we pursue an iterative improvement scheme to extend the admissible range of $\gamma$ up to the optimal value $\gamma=1$.
By employing the decay rate for the coefficient $\frac{\varphi}{\tau}\lesssim e^{-T}$, and applying the estimate (4.1), setting $\gamma=\frac{1}{2}$ in (159) yields
(160) $e^{T}\mathscr{E}^{low}(T)\lesssim 1+(\mathcal{I}^{0})^{2}+\mathcal{O}^{2}.$
Using this improved decay for the energy, we revisit the energy inequality. Incorporating the refined decay estimate for $\nabla^{I}\Sigma||_{L^{2}(M)}$,
$\nabla^{I}\Sigma||_{L^{2}(M)}\lesssim e^{-\frac{1}{2}T}\left(\mathcal{O}+(\mathcal{I}^{0})^{2}+1\right),\quad I\leq 2,$
corresponding to $\gamma=\frac{1}{2}$, the differential inequality governing $\mathscr{E}^{low}(T)$ takes the form
$\frac{d}{dT}\mathscr{E}^{low}+2(n-1)\mathscr{E}^{low}\lesssim e^{-(1+\frac{1}{2})T}\mathcal{O}^{2}+e^{-(1+4\gamma)T}\mathds{Y}^{4}+e^{-(1+3\gamma)T}\mathds{Y}^{3}+e^{-4\gamma T}\mathds{Y}^{4}+e^{-(1+3\gamma)T}\mathcal{I}^{0}\mathds{Y}^{3}.$
Integrating this inequality, one obtains the improved bound
(161) $e^{(1+\frac{1}{2})T}\mathscr{E}^{low}(T)\lesssim 1+(\mathcal{I}^{0})^{2}+\mathcal{O}^{2}.$
Iterating this procedure—each time replacing $\gamma$ by the newly achieved decay rate—leads to a convergent sequence for $2\gamma^{optimal}$ given by
(162) $2\gamma^{optimal}=1+\frac{1}{2}\left(1+\frac{1}{2}\left(1+\cdots\right)\right)=2,\quad\text{hence}\quad\gamma^{optimal}=1.$
Consequently, the optimal decay rate for the low-order energy is established:
(163) $e^{2\gamma T}\mathscr{E}^{low}(T)\lesssim 1+(\mathcal{I}^{0})^{2}+\mathcal{O}^{2},\quad\gamma=1.$
As an immediate corollary, we obtain
Corollary 4.1.
$\mathcal{F}\lesssim 1+\mathcal{I}^{0}+\mathcal{O}$ and $\mathcal{N}^{\infty}\lesssim 1+(\mathcal{I}^{0})^{2}+\mathcal{O}^{2}$ by elliptic estimates 4.1.
4.3. Energy Estimates
With the individual estimates available, we are ready to complete the energy estimates for the pair $(\Sigma,\mathfrak{T})$. Let us recall the definition of the top order energy $\mathscr{E}^{top}$
Definition 7.
(164) $\mathscr{E}^{top}:=\frac{1}{2}\sum_{I\leq 3}\int_{M}\langle\nabla^{I}\Sigma,\nabla^{I}\Sigma\rangle\mu_{g}+\frac{1}{2}\sum_{I\leq 3}\int_{M}\langle\nabla^{I-1}\mathfrak{T},\nabla^{I-1}\mathfrak{T}\rangle\mu_{g}.$
First, we recall the following elementary proposition
Proposition 4.2 (Integrated energy estimate for the wave system).
Let $(M,g)$ be a closed Riemannian manifold and consider a time interval $T_{0},T_{\infty}]\subset\mathbb{R}$. Suppose $(\Phi,\Psi)$ is a pair of smooth tensor fields of type $(K,L)$ on $T_{0},T_{\infty}]\times M$ satisfying the coupled system
(165) $\partial_{T}\Phi$ $=-N\Psi+P_{\Phi},$
(166) $\partial_{T}\Psi$ $=-N\Delta_{g}\Phi+Q_{\Psi},$
where $N=N(T,x)$ is the lapse function, and $P_{\Phi}$, $Q_{\Psi}$ are given source terms of the same tensorial type as $\Phi$ and $\Psi$, respectively. Denote by $M(T)$ the time slice at time $T$ equipped with the induced metric $g(T)$ and volume form $\mu_{g}$.
Then, the following integrated energy identity holds for all $T\in[T_{0},T_{\infty}]$:
(167) $\int_{M(T)}\big(|\nabla\Phi|^{2}+|\Psi|^{2}\big)\,\mu_{g}=$ $\int_{M(T_{0})}\big(|\nabla\Phi|^{2}+|\Psi|^{2}\big)\,\mu_{g}$
$+2\int_{T_{0}}^{T}\int_{M(t)}\big(\langle\nabla P_{\Phi},\nabla\Phi\rangle+\langle Q_{\Psi},\Psi\rangle\big)\,\mu_{g}\,dt$
$+\int_{T_{0}}^{T}\mathscr{E}\mathscr{R}(t)\,dt,$
where the error functional $\mathscr{E}\mathscr{R}(t)$ admits the bound
(168) $|\mathscr{E}\mathscr{R}(t)|\lesssim$ $\left\|\frac{N}{n}-1\right\|_{L^{\infty}(M(t))}\big(\|\nabla\Phi\|_{L^{2}(M(t))}^{2}+\|\Psi\|_{L^{2}(M(t))}^{2}\big)$
$+\frac{\varphi}{\tau}\|\Sigma\|_{L^{\infty}(M(t))}\big(\|\nabla\Phi\|_{L^{2}(M(t))}^{2}+\|\Psi\|_{L^{2}(M(t))}^{2}\big)$
$+\left[\frac{\varphi}{\tau}\big(\|\nabla\Sigma\|_{L^{4}(M(t))}+\|\Sigma\|_{L^{\infty}(M(t))}\|\nabla(\frac{N}{n}-1)\|_{L^{4}(M(t))}\big)+\|\nabla(\frac{N}{n}-1)\|_{L^{4}(M(t))}\right]$
$\quad\times\|\Phi\|_{L^{4}(M(t))}\|\nabla\Phi\|_{L^{2}(M(t))}.$
Here, all norms are computed with respect to the metric $g(T)$ on the slice $M(T)$, and the constants implied by $\lesssim$ are of numerical type.
Proof.
Let $(\Phi,\Psi)$ be a smooth pair of tensor fields on the compact Riemannian manifold $(M,g)$, where the metric $g=g(T)$ depends smoothly on time $T\in[T_{0},T_{\infty}]$. We consider the energy functional
$\mathscr{E}(T):=\frac{1}{2}\int_{M(T)}\big(|\nabla\Phi|^{2}+|\Psi|^{2}\big)\,\mu_{g},$
where $\mu_{g}$ denotes the Riemannian volume form induced by $g(T)$ on the hypersurface $M(T)$, and $|\cdot|$ and $\nabla$ denote the metric norm and Levi-Civita connection respectively.
By the transport theorem for evolving hypersurfaces (cf. [32]), we have the exact identity
(169) $\frac{d}{dT}\int_{M(T)}\left(|\nabla\Phi|^{2}+|\Psi|^{2}\right)\mu_{g}=\int_{M(T)}\partial_{T}\left(|\nabla\Phi|^{2}+|\Psi|^{2}\right)\mu_{g}+\int_{M(T)}\left(|\nabla\Phi|^{2}+|\Psi|^{2}\right)\partial_{T}\mu_{g}.$
We rewrite (169) as
(170) $\frac{d}{dT}\mathscr{E}(T)=\int_{M(T)}\left(\langle\nabla\partial_{T}\Phi,\nabla\Phi\rangle+\langle\partial_{T}\Psi,\Psi\rangle\right)\mu_{g}+\mathscr{ER}(T),$
where the error term $\mathscr{ER}(T)$ collects the contributions
$\mathscr{ER}(T)\sim\int_{M(T)}\langle[\partial_{T},\nabla]\Phi,\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\partial_{T}(g^{-1}....g^{-1}g.....g)\nabla\Phi\nabla\Phi\mu_{g}$
$\int_{M(T)}\partial_{T}(g^{-1}....g^{-1}g.....g)\Psi\Psi\mu_{g}+\int_{M(T)}(|\nabla\Phi|^{2}+|\Psi|^{2})\partial_{T}\mu_{g}$
Recall that the time evolution of the metric $g=g(T)$
(171) $\partial_{T}g_{ij}=-\frac{2\varphi}{\tau}N\Sigma_{ij}-2\left(1-\frac{N}{n}\right)g_{ij},\quad\text{and}\quad\partial_{T}g^{ij}=\frac{2\varphi}{\tau}N\Sigma^{ij}+2\left(1-\frac{N}{n}\right)g^{ij},$
where $\varphi,\tau,N,n$, and $\Sigma$ are as in the geometric setup.
Consequently, the time derivative of the volume form $\mu_{g}$ is given by
$\partial_{T}\mu_{g}=\frac{1}{2}\mu_{g}\,\mathrm{tr}_{g}(\partial_{T}g)=-n\left(1-\frac{N}{n}\right)\mu_{g},$
which follows from (171).
The commutator $\partial_{T},\nabla]\Phi$ is expressible via the time derivative of the Christoffel symbols $\Gamma[g]$:
$\partial_{T},\nabla]\Phi=(\partial_{T}\Gamma)\ast\Phi,$
where $\ast$ denotes appropriate tensor contractions. By direct computation,
$|\partial_{T}\Gamma|\lesssim\frac{\varphi}{\tau}|\nabla(N\Sigma)|+|\nabla\left(\tfrac{N}{n}-1\right)|,$
hence
$|[\partial_{T},\nabla]\Phi|\lesssim\frac{\varphi}{\tau}\left(|\nabla\Sigma|+|\Sigma||\nabla(\tfrac{N}{n}-1)|\right)+|\nabla(\tfrac{N}{n}-1)|.$
Applying Hölder’s inequality with Sobolev embeddings yields
$\left|\int_{M(T)}\langle[\partial_{T},\nabla]\Phi,\nabla\Phi\rangle\mu_{g}\right|$ $\lesssim\frac{\varphi}{\tau}\left(\|\nabla\Sigma\|_{L^{4}}+\|\Sigma\|_{L^{\infty}}\|\nabla(\tfrac{N}{n}-1)\|_{L^{4}}\right)\|\Phi\|_{L^{4}}\|\nabla\Phi\|_{L^{2}}$
$\quad+\|\nabla(\tfrac{N}{n}-1)\|_{L^{4}}\|\Phi\|_{L^{4}}\|\nabla\Phi\|_{L^{2}}.$
Similarly, the terms involving $\partial_{T}\mu_{g}$ are controlled by
$\left|\int_{M(T)}\left(|\nabla\Phi|^{2}+|\Psi|^{2}\right)\partial_{T}\mu_{g}\right|\lesssim\left\|1-\frac{N}{n}\right\|_{L^{\infty}}\left(\|\nabla\Phi\|_{L^{2}}^{2}+\|\Psi\|_{L^{2}}^{2}\right).$
Collecting these bounds, we deduce the estimate
(172) $|\mathscr{ER}(T)|\lesssim$ $\left(\left\|\frac{N}{n}-1\right\|_{L^{\infty}(M(T))}+\frac{\varphi}{\tau}\|\Sigma\|_{L^{\infty}(M(T))}\right)\left(\|\nabla\Phi\|_{L^{2}(M(T))}^{2}+\|\Psi\|_{L^{2}(M(T))}^{2}\right)$
$+\left[\frac{\varphi}{\tau}\left(\|\nabla\Sigma\|_{L^{4}(M(T))}+\|\Sigma\|_{L^{\infty}(M(T))}\|\nabla(\tfrac{N}{n}-1)\|_{L^{4}(M(T))}\right)+\|\nabla(\tfrac{N}{n}-1)\|_{L^{4}(M(T))}\right]\|\Phi\|_{L^{4}(M(T))}\|\nabla\Phi\|_{L^{2}(M(T))}.$
We now turn to the principal terms in (170):
$\int_{M(T)}\langle\nabla\partial_{T}\Phi,\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\langle\partial_{T}\Psi,\Psi\rangle\mu_{g}.$
Substituting the evolution equations
$\partial_{T}\Phi=-N\Psi+P_{\Phi},\quad\partial_{T}\Psi=-N\Delta_{g}\Phi+Q_{\Psi},$
we write
$\int_{M(T)}\langle\nabla\partial_{T}\Phi,\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\langle\partial_{T}\Psi,\Psi\rangle\mu_{g}$
$=\int_{M(T)}\langle\nabla(-N\Psi+P_{\Phi}),\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\langle-N\Delta_{g}\Phi+Q_{\Psi},\Psi\rangle\mu_{g}$
$=\int_{M(T)}\langle\nabla P_{\Phi},\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\langle Q_{\Psi},\Psi\rangle\mu_{g}+\underbrace{\int_{M(T)}\langle\nabla(-N\Psi),\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\langle-N\Delta_{g}\Phi,\Psi\rangle\mu_{g}}_{=:I}.$
By integration by parts and the absence of boundary (since $M$ is closed), the integral
$I=\int_{M(T)}\langle\nabla(-N\Psi),\nabla\Phi\rangle\mu_{g}+\int_{M(T)}\langle-N\Delta_{g}\Phi,\Psi\rangle\mu_{g}$
vanishes exactly. This cancellation is a fundamental feature reflecting the underlying hyperbolic structure of the system and guarantees no loss of derivatives in the energy estimates, preserving regularity. This completes the proof. ∎
Remark 10.
[…]
Theorem 4.2 (Top-order energy bound).
Assume the hypotheses of the main theorem 1.1 are satisfied, and let $(\Sigma,\mathfrak{T})$ denote the coupled wave pair associated with the rescaled Einstein system in the gauge fixed framework of this work. Let $\mathscr{E}^{\mathrm{top}}(T)\geq 0$ denote the top–order energy functional corresponding to $(\Sigma,\mathfrak{T})$ as defined in 7, and let $\mathcal{I}^{0}$ denote the initial size of the data at time $T_{0}=a\gg 1$.
Then for every $T$ in the lifespan of the solution one has
$\mathscr{E}^{\mathrm{top}}\ \lesssim\ 1+\mathcal{I}^{0},$
where the involved constant is of purely numerical type.
Proof.
The proof of the energy estimates for the top-order term is based on the proposition 4.2. First, recall the definition of the top-order energy
(173) $\mathscr{E}^{top}:=\frac{1}{2}\sum_{I\leq 2}\int_{M}\langle\nabla^{I+1}\Sigma,\nabla^{I+1}\Sigma\rangle\mu_{g}+\frac{1}{2}\sum_{I\leq 2}\int_{M}\langle\nabla^{I}\mathfrak{T},\nabla^{I}\mathfrak{T}\rangle\mu_{g}.$
We want to use the proposition 4.2 with $\Phi=\nabla^{I}\Sigma$ and $\Psi=\nabla^{I}\mathfrak{T}$ i.e.,
(174) $\partial_{T}\nabla^{I}\Sigma=-N\frac{\varphi}{\tau}\nabla^{I}\mathfrak{T}+P^{I},$
(175) $\partial_{T}\nabla^{I}\mathfrak{T}=-N\frac{\varphi}{\tau}\Delta_{g}\nabla^{I}\Sigma+Q^{I},$
where the error terms $P^{I}$ are $Q^{I}$ are expressed schematically as follows
$P^{I}$ $:=$ $[\partial_{T},\nabla^{I}]\Sigma+\frac{\varphi}{\tau}[\nabla^{I},N]\mathfrak{T}+\nabla^{I}\Bigg(-(n-1)\Sigma_{ij}+\frac{\varphi}{\tau}\nabla_{i}\nabla_{j}\left(\frac{N}{n}-1\right)$
(176) $\quad+\frac{2\varphi}{\tau}N\Sigma_{ik}\Sigma^{k}_{j}-\frac{\varphi}{n\tau}\left(\frac{N}{n}-1\right)g_{ij}-(n-2)\left(\frac{N}{n}-1\right)\Sigma_{ij}\Bigg),$
$Q^{I}$ $:=$ $[\partial_{T},\nabla^{I}]\mathfrak{T}-\frac{\varphi}{\tau}[\nabla^{I},N\Delta_{g}]\Sigma+\nabla^{I}\Bigg(n\frac{\varphi}{\tau}\nabla^{l}\nabla_{i}\left(\frac{N}{n}-1\right)\Sigma_{jl}$
$\quad+n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{i}\Sigma_{jl}+(2-n)\nabla_{i}\nabla_{j}\left(\frac{N}{n}-1\right)$
$\quad+n\frac{\varphi}{\tau}\nabla^{l}\nabla_{j}\left(\frac{N}{n}-1\right)\Sigma_{il}+n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{j}\Sigma_{il}$
$\quad-n\frac{\varphi}{\tau}\Delta_{g}\left(\frac{N}{n}-1\right)\Sigma_{ij}-2n\frac{\varphi}{\tau}\nabla^{l}\left(\frac{N}{n}-1\right)\nabla_{l}\Sigma_{ij}$
(177) $\quad-\Delta_{g}\left(\frac{N}{n}-1\right)g_{ij}+N\frac{\varphi}{\tau}\left(\mathfrak{T}_{ki}\Sigma^{k}_{j}+\mathfrak{T}_{kj}\Sigma^{k}_{i}\right)+\frac{2(n-1)}{n^{2}}\left(\frac{N}{n}-1\right)g_{ij}\Bigg).$
A couple of important points to note here. In the expression or $P^{I}$, term $\frac{\varphi}{n\tau}\nabla^{I}(\frac{N}{n}-1)g_{ij}$ is pure trace while $\Sigma$ is transe-verse traceless. Therefore, this term does not contribute to the energy estimates. Similarly, in the expression of $Q^{I}$, the terms $\nabla^{I+2}(\frac{N}{n}-1)g_{ij}$ and $\nabla^{I}(\frac{N}{n}-1)g_{ij}$ are of pure trace type and therefore in the energy estimate for $\mathfrak{T}$ they contribute to the nonlinear term $|\Sigma|^{2}$ since $\mbox{tr}_{g}\mathfrak{T}=|\Sigma|^{2}$ by the Hamiltonian constraint (67). Now apply proposition 4.2 to the system (174)-(175) to obtain
$\mathscr{E}^{top}(T)=\mathscr{E}^{top}(T_{0})+2\sum_{I\leq 2}\int_{T_{0}}^{T}\int_{M(T)}\left(\langle\nabla P^{I},\nabla^{I+1}\Sigma\rangle+\langle Q^{I},\nabla^{I}\mathfrak{T}\rangle\right)\mu_{g}dt+\int_{T_{0}}^{T}\mathscr{ER}(t)dt,$
where
[…]
Our goal is to control the error terms and the spacetime integral terms involving $P^{I}$ and $Q^{I}$. We do so schematically as follows
(178) $\int_{T_{0}}^{T}\int_{M(T)}\langle\nabla P^{I},\nabla^{I+1}\Sigma\rangle\mu_{g}dt$
$=\underbrace{-2(n-1)\int_{T_{0}}^{T}\int_{M(T)}|\nabla^{I+1}\Sigma|^{2}\mu_{g}dt}_{I-decay~term}$
$+\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\Sigma\nabla^{I+1}\Sigma\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}+J_{3}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}+1}\Sigma\nabla^{J_{3}+I-m-1}\Sigma\nabla^{I+1}\Sigma\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\Sigma\nabla^{I+1}\Sigma\mu_{g}dt$
$+\sum_{I\leq 2}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{I+3}(\frac{N}{n}-1)\nabla^{I+1}\Sigma\mu_{g}$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}+J_{3}=I+1}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}}\Sigma\nabla^{J_{3}}\Sigma\nabla^{I+1}\Sigma\mu_{g}$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I+1}\int_{T_{0}}^{T}\int_{M(T)}\nabla^{J_{1}}(\frac{N}{n}-1)\nabla^{J_{2}}\Sigma\nabla^{I+1}\Sigma\mu_{g}$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}}\mathfrak{T}\nabla^{I+1}\Sigma\mu_{g}.$
We can estimate each term as follows. We use the elliptic estimate whenever necessary. First note that $\frac{\varphi}{\tau}=\frac{e^{-T}}{\sqrt{e^{-2T}+3n(n+1)}}\lesssim e^{-T}$
$|\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m}\Sigma\nabla^{I+1}\Sigma|\lesssim\int_{T_{0}}^{T}e^{-(1+3\gamma)t}\mathds{Y}^{4}dt$
$\lesssim(e^{-(1+3\gamma)T_{0}}-e^{-(1+3\gamma)T})\mathds{Y}^{3}\Gamma\lesssim e^{-(1+3\gamma)T_{0}}(1-e^{-(1+3\gamma)(T-T_{0})})\mathds{Y}^{3}\Gamma\lesssim 1.$
The maximum derivative on lapse function in the previous estimate is $I+1$ and therefore by elliptic estimate, it is bounded by $\Sigma||^{2}_{H^{I-1}(M)}$ which yields $e^{-2\gamma t}$ decay. The next term is estimated as
$|\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}+J_{3}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}+1}\Sigma\nabla^{J_{3}+I-m}\Sigma\nabla^{I+1}\Sigma\mu_{g}dt|\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\mathds{Y}\Gamma^{2}dt$
$\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\mathds{Y}\Gamma^{2}\lesssim 1.$
The next term is estimated as
(179) $|\sum_{I\leq 2}\sum_{m=0}^{I}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m}\Sigma\nabla^{I+1}\Sigma\mu_{g}dt|\lesssim\int_{T_{0}}^{T}e^{-3\gamma t}\Gamma\mathds{Y}^{3}dt$
$\lesssim e^{-3\gamma T_{0}}(1-e^{-3\gamma(T-T_{0})})\Gamma\mathds{Y}^{3}\lesssim 1,$
(180) $|\sum_{I\leq 2}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{I+3}(\frac{N}{n}-1)\nabla^{I+1}\Sigma\mu_{g}|\lesssim\int_{T_{0}}^{T}e^{-t}\Gamma^{3}dt\lesssim e^{-T_{0}}(1-e^{-(T-T_{0})})\Gamma^{3}\lesssim 1,$
(181) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}+J_{3}=I+1}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}}\Sigma\nabla^{J_{3}}\Sigma\nabla^{I+1}\Sigma\mu_{g}|\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\Gamma^{2}\mathds{Y}dt$
$\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}\lesssim 1.$
The next term is estimated as follows
(182) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I+1}\int_{T_{0}}^{T}\int_{M(T)}\nabla^{J_{1}}(\frac{N}{n}-1)\nabla^{J_{2}}\Sigma\nabla^{I+1}\Sigma\mu_{g}|\lesssim\int_{T_{0}}^{T}e^{-2\gamma t}\Gamma^{2}\mathds{Y}^{2}dt$
$\lesssim e^{-2\gamma T_{0}}(1-e^{-2\gamma(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1,$
where we noted that the top order term in lapse $\nabla^{4}(\frac{N}{n}-1)||_{L^{2}(M)}$ is estimated by $\Sigma||^{2}_{H^{2}(M)}$ which exhibits $e^{-2\gamma t}$ decay. The last term in the expression of $P^{I}$ is estimated as
(183) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}}\mathfrak{T}\nabla^{I+1}\Sigma\mu_{g}|\lesssim\int_{T_{0}}^{T}e^{-(1+2\gamma)t}\Gamma^{2}\mathds{Y}^{2}dt$
$\lesssim e^{-(1+2\gamma)T_{0}}(1-e^{-(1+2\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1.$
Now we estimate the terms in the expressions of $Q^{I}$. First, the spacetime integral involving $Q^{I}$ is explicitly evaluated as follows (written in schematic notation)
(184) $\int_{T_{0}}^{T}\int_{M(T)}\langle Q^{I},\nabla^{I}\mathfrak{T}\rangle\mu_{g}dt$
$\sim\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\mathfrak{T}\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}+J_{3}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}+1}\Sigma\nabla^{J_{3}+I-m-1}\mathfrak{T}\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\mathfrak{T}\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M}\nabla(\frac{N}{n}-1)\nabla^{I+1}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}N\nabla^{J_{1}}\text{Riem}\nabla^{J_{2}+I-m}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}\text{Riem}\nabla^{J_{2}+I-m-1}\Sigma,\nabla^{I}\mathfrak{T}\rangle\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}\text{Riem}\nabla^{J_{2}}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+2}(\frac{N}{n}-1)\nabla^{J_{2}}\Sigma,\nabla^{I}\mathfrak{T}\rangle\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+1}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M}\langle\nabla^{J_{1}}\mathfrak{T}\nabla^{J_{2}}\Sigma\nabla^{I}\mathfrak{T}\rangle\mu_{g}dt$
$+\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I-1}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+1}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt.$
Now we estimate each term separately. The first term is estimated as follows
(185) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\mathfrak{T}\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+2\gamma)t}\mathds{Y}^{2}\Gamma^{2}dt\lesssim e^{-(1+2\gamma)T_{0}}(1-e^{-(1+2\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1.$
Here, the elliptic estimate 4.1 for the lapse function is used. The next term is estimated as
(186) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}+J_{3}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}N\nabla^{J_{2}+1}\Sigma\nabla^{J_{3}+I-m-1}\mathfrak{T}\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\Gamma^{2}\mathds{Y}dt\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}\lesssim 1.$
Notice that $\Sigma$ appears as $\nabla^{I}\Sigma$ at the top most order and therefore is estimated in $L^{2}$ yielding extra $e^{-\gamma t}$ decay. The next term is estimated as
(187) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+I-m-1}\mathfrak{T}\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-2\gamma t}\mathds{Y}^{2}\Gamma^{2}dt\lesssim e^{-2\gamma T_{0}}(1-e^{-2\gamma(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1.$
The next terms are estimated as
(188) $|\sum_{I\leq 2}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M}\nabla(\frac{N}{n}-1)\nabla^{I+1}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|\lesssim\int_{T_{0}}^{T}e^{-(1+2\gamma)t}\mathds{Y}^{2}\Gamma^{2}dt$
$\lesssim e^{-(1+2\gamma)T_{0}}(1-e^{-(1+2\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1,$
(189) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}N\nabla^{J_{1}}\text{Riem}\nabla^{J_{2}+I-m}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\Gamma^{2}\mathds{Y}dt\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}\lesssim 1,$
(190) $|\sum_{I\leq 2}\sum_{m=0}^{I-1}\sum_{J_{1}+J_{2}=m}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}\text{Riem}\nabla^{J_{2}+I-m-1}\Sigma,\nabla^{I}\mathfrak{T}\rangle\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\Gamma^{2}\mathds{Y}dt\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}\lesssim 1,$
(191) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}}\text{Riem}\nabla^{J_{2}}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\Gamma^{2}\mathds{Y}dt\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}\lesssim 1,$
(192) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+2}(\frac{N}{n}-1)\nabla^{J_{2}}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+3\gamma)t}\mathds{Y}^{3}\Gamma dt\lesssim e^{-(1+3\gamma)T_{0}}(1-e^{-(1+3\gamma)(T-T_{0})})\Gamma\mathds{Y}^{3}\lesssim 1,$
(193) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+1}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+2\gamma)t}\mathds{Y}^{2}\Gamma^{2}dt\lesssim e^{-(1+2\gamma)T_{0}}(1-e^{-(1+2\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1,$
(194) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M}\nabla^{J_{1}}\mathfrak{T}\nabla^{J_{2}}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|\lesssim\int_{T_{0}}^{T}e^{-(1+\gamma)t}\mathds{Y}\Gamma^{2}dt$
$\lesssim e^{-(1+\gamma)T_{0}}(1-e^{-(1+\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}\lesssim 1,$
and
(195) $|\sum_{I\leq 2}\sum_{J_{1}+J_{2}=I-1}\int_{T_{0}}^{T}\frac{\varphi}{\tau}\int_{M(T)}\nabla^{J_{1}+1}(\frac{N}{n}-1)\nabla^{J_{2}+1}\Sigma\nabla^{I}\mathfrak{T}\mu_{g}dt|$
$\lesssim\int_{T_{0}}^{T}e^{-(1+2\gamma)t}\Gamma^{2}\mathds{Y}^{2}dt\lesssim e^{-(1+2\gamma)T_{0}}(1-e^{-(1+2\gamma)(T-T_{0})})\Gamma^{2}\mathds{Y}^{2}\lesssim 1.$
Now note that the first term in the expression involving $P^{I}$ contains term $I$ which is negative definite. Therefore, the collection of every term yields
(196) $\mathscr{E}^{top}(T)\lesssim\mathscr{E}^{top}(T_{0})+1\lesssim\mathcal{I}^{0}+1$
which completes the proof of the theorem. ∎
Corollary 4.2.
$\mathcal{O}\lesssim 1+\mathcal{I}^{0}$
Remark 11.
Notice that the estimates are uniform in $T$, so one can take the limit $T_{\infty}=\infty$.
5. Proof of the main results
5.1. Proof of theorem 1.1
In this section, we prove the main theorem with the aid of the uniform estimates obtained in section 4.3. Note that we have already presented the idea of the proof in section 3.3. Here we will use a contradiction argument. First, recall the statement of the main theorem
Theorem (Global existence, $\Lambda>0,\sigma(M)\leq 0$).
[…]
Fix a smooth background Riemannian metric $\xi_{0}$ on $M$ and a constant $C>1$. For any sufficiently large initial energy quantity $\mathcal{I}^{0}>0$, there exists a constant $a=a(\mathcal{I}^{0})>0$, sufficiently large so that $\mathcal{I}^{0}e^{-a/10}<1$.
Let $(g_{0},\Sigma_{0})$ be an initial data set verifying the Einstein constraint equations at initial CMC time $T_{0}=a$, written in CMC-transported spatial coordinates and satisfying:
(197) $C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},$
(198) $\sum_{I=0}^{3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{I=0}^{2}\left(\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}+\|e^{a}\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\right)\leq\mathcal{I}^{0},$
where $\mathfrak{T}[g_{0}]$ denotes the renormalized trace-free spatial Ricci curvature tensor of $g_{0}$.
Then, the Einstein-$\Lambda$ evolution equations admit a unique classical solution
$T\mapsto(g(T),\Sigma(T))\in\mathcal{C}^{\infty}([T_{0},\infty)\times M)$
in CMC-transported spatial coordinates, satisfying the constraint equations at each slice $T$ and obeying the following uniform a priori estimates for all $T\in[T_{0},\infty)$:
(199) $\sum_{I=0}^{3}\|\nabla^{I}\Sigma(T)\|_{L^{2}(M)}+\sum_{I=0}^{2}\left(\|\nabla^{I}\mathfrak{T}[g(T)]\|_{L^{2}(M)}+\|e^{T}\nabla^{I}\Sigma(T)\|_{L^{2}(M)}\right)\leq C_{1}(1+\mathcal{I}^{0}),$
(200) $C_{2}^{-1}g_{0}\leq g(T)\leq C_{2}g_{0}.$
Here, $C_{1},C_{2}>0$ are numerical constants depending only on the universal geometric and analytic data of the problem (e.g., Sobolev constants of $(M,g_{0})$ and the constants in the structure equations), but independent of $T$. The developed spacetime is future geodesically complete.
Moreover, the solution $(g(T),\Sigma(T))$ converges in the $C^{\infty}$ topology, as $T\to\infty$, to a limiting Riemannian metric $\widetilde{g}$ on MM of pointwise constant negative scalar curvature, in the sense that
$\Sigma(T)\to 0\quad\text{and}\quad g(T)\to\widetilde{g},\quad\text{as }T\to\infty,$
[…]
We begin by establishing global existence of the CMC Einstein-$\Lambda$ flow under the stated assumptions. The proof proceeds by a standard continuation argument, combining local well-posedness for the elliptic-hyperbolic system with a contradiction argument based on a priori estimates and weak compactness.
Let us first observe that the coupled Einstein-$\Lambda$ system, expressed in CMC-transported spatial coordinates, reduces to a manifestly elliptic-hyperbolic formulation for the unknowns $(g,\Sigma)$, consisting of a hyperbolic evolution equation for the metric $g$, a transport-type equation for $\Sigma$, and elliptic constraint equations for the lapse and shift. The system satisfies the structural conditions of the classical theory developed for mixed elliptic-hyperbolic systems (see, e.g., [11]). In particular, the associated initial value problem admits a unique classical solution on a time interval $T_{0},T)$ for some $T>T_{0}$, with the initial data $(g_{0},\Sigma_{0})$ prescribed at time $T_{0}=a$.
We assume, for contradiction, that the maximal existence interval is bounded above by some finite time $T<\infty$, i.e., the solution ceases to exist beyond time $T$. By the uniform a priori estimates derived in Section 4.3, combined with Sobolev persistence of regularity, the fields $g(T^{\prime}),\Sigma(T^{\prime}),\mathfrak{T}(T^{\prime})$ remain uniformly bounded in the Sobolev spaces $H^{3}(M)$, $H^{3}(M)$, and $H^{2}(M)$ respectively, for all $T^{\prime}\in[T_{0},T)$.
Using reflexivity and the Banach-Alaoglu theorem, we extract a weakly convergent subsequence $\{T_{n}^{\prime}\}_{n\in\mathbb{N}}\subset[T_{0},T)$ with $T_{n}^{\prime}\to T^{-}$ such that:
$\Sigma(T_{n}^{\prime})$ $\rightharpoonup\Sigma(T)\quad\text{weakly in }H^{3}(M),$
$\mathfrak{T}(T_{n}^{\prime})$ $\rightharpoonup\mathfrak{T}(T)\quad\text{weakly in }H^{2}(M).$
Moreover, by weak lower semicontinuity of the Sobolev norms and the uniform estimates, we have:
$\sum_{I=0}^{3}\|\nabla^{I}\Sigma(T)\|_{L^{2}(M)}$ $\leq\liminf_{n\to\infty}\sum_{I=0}^{3}\|\nabla^{I}\Sigma(T_{n}^{\prime})\|_{L^{2}(M)}\lesssim 1+\mathcal{I}^{0},$
$\sum_{I=0}^{2}\|\nabla^{I}\mathfrak{T}(T)\|_{L^{2}(M)}$ $\leq\liminf_{n\to\infty}\sum_{I=0}^{2}\|\nabla^{I}\mathfrak{T}(T_{n}^{\prime})\|_{L^{2}(M)}\lesssim 1+\mathcal{I}^{0}.$
Since the constraint equations are preserved along the flow and depend smoothly on the variables, it follows that $(g(T),\Sigma(T))$ defines a compatible initial data set at time $T$. By the same local well-posedness theory, we may extend the solution to a strictly larger time interval $T,T+\epsilon)$ for some $\epsilon>0$ depending only on the bounds for $\Sigma(T)$ and $\mathfrak{T}(T)$. This contradicts the assumption that $T_{0},T)$ was the maximal interval of existence. We therefore conclude that the solution extends globally in time, i.e., $T=\infty$.
We now address the asymptotic behavior as $T\to\infty$. From the a priori energy estimates derived in Section 4.3, and in particular from exponential decay of the appropriate norms, the time derivative of the metric satisfies:
$\|\partial_{T}g(T^{\prime})\|_{H^{2}(M)}\lesssim\mathcal{I}^{0}e^{-(1+\gamma)T^{\prime}}+(\mathcal{I}^{0})^{2}e^{-2\gamma T^{\prime}}$
for all $T^{\prime}\geq T_{0}$. Integrating this differential inequality from $T$ to infinity yields
$\|g(T)-g(\infty)\|_{H^{2}(M)}$ $\leq\int_{T}^{\infty}\|\partial_{T^{\prime}}g(T^{\prime})\|_{H^{3}(M)}\,dT^{\prime}$
$\lesssim\mathcal{I}^{0}e^{-(1+\gamma)T}+(\mathcal{I}^{0})^{2}e^{-2\gamma T},$
which shows that $g(T)$ converges strongly in $H^{2}(M)$ to a limiting Riemannian metric $\widetilde{g}:=g(\infty)$ as $T\to\infty$. Similarly,
$\|\Sigma(T)\|_{H^{2}(M)}\lesssim\mathcal{I}^{0}e^{-\gamma T}\to 0,$
so that $\Sigma(T)\to 0$ strongly in $H^{2}(M)$, and hence in $C^{0}(M)$ by Sobolev embedding.
Finally, we consider the scalar curvature of the limiting metric. The Hamiltonian constraint equation takes the form
$R(g(T))+\frac{n-1}{n}=|\Sigma(T)|^{2}_{g},$
and from the above convergence and the Sobolev algebra property of $H^{s}(M)$ for $s>\frac{3}{2}$, it follows that:
$\|R(g(T))+\tfrac{n-1}{n}\|_{H^{2}(M)}$ $=\||\Sigma(T)|^{2}_{g}\|_{H^{2}(M)}$
$\lesssim\|\Sigma(T)\|_{H^{2}(M)}^{2}\to 0.$
Therefore, in the limit $T\to\infty$, we obtain:
$R(\widetilde{g})=\lim_{T\to\infty}R(g(T))=-\frac{n-1}{n},$
[…]
$\lambda(T)\to\infty\quad\text{as}\quad T\to\infty\quad\Longleftrightarrow\quad\int_{T_{0}}^{\infty}\frac{1}{\alpha^{0}(T)}\,dT=\infty.$
Let $N$ denote the rescaled lapse in the CMC gauge so that the spacetime metric reads
$\widehat{g}=-N^{2}dT^{2}+g_{ij}(T)dx^{i}dx^{j}.$
Introduce the future-directed co-vector field $Z:=N\partial_{T}$, and decompose the geodesic tangent as
$\alpha=N\alpha^{0}Z+W,$
where $W$ is tangent to the constant-$T$ hypersurfaces. Then
$\widehat{g}(\alpha,\alpha)=-N^{2}(\alpha^{0})^{2}+|W|^{2}_{g(T)}=\begin{cases}-1&\text{(timelike)}\\ 0&\text{(null)}\end{cases},$
so that $|W|^{2}_{g(T)}=N^{2}(\alpha^{0})^{2}+\epsilon$, where $\epsilon=1$ or 0 respectively. Thus,
$|W|^{2}_{g(T)}\geq N^{2}(\alpha^{0})^{2},\quad\text{and in all cases}\quad f:=N^{2}(\alpha^{0})^{2}\geq 0.$
Differentiating $f$ along $\mathcal{C}$ and using the geodesic equation $\nabla[\widehat{g}]_{\alpha}\alpha=0$, we obtain
$\frac{d}{dT}f=\frac{2}{\alpha^{0}}\widehat{g}(\alpha,Z)\cdot\widehat{g}(\alpha,\nabla[\widehat{g}]_{\alpha}Z),$
and using $\widehat{g}(\alpha,Z)=-N\alpha^{0}$, we compute
$\widehat{g}(\alpha,\nabla_{\alpha}Z)=-\alpha^{0}\nabla_{W}N+K_{ij}W^{i}W^{j},$
where $K$ is the second fundamental form of the constant-$T$ slices. Thus,
$\frac{d}{dT}f=-2N\left(\alpha^{0}\nabla_{W}N-K_{ij}W^{i}W^{j}\right).$
Decompose $K=\Sigma+\tfrac{\tau}{n}g$, with $\tau=\mathrm{tr}_{g}K<0$ and $\Sigma$ trace-free. Using Cauchy–Schwarz and Sobolev embedding, we estimate
$\left|\frac{d}{dT}\log f\right|\leq\|\nabla N\|_{L^{\infty}}+\|N\Sigma\|_{L^{\infty}}.$
By the main decay estimates (Theorem 1.1), we have
$\|\nabla N\|_{L^{\infty}}+\|N\Sigma\|_{L^{\infty}}\lesssim e^{-T},$
and thus
$\left|\frac{d}{dT}\log f\right|\lesssim e^{-T},\quad\Rightarrow\quad f(T)\to C>0\quad\text{as }T\to\infty.$
Therefore, $N^{2}(\alpha^{0})^{2}\leq C$ uniformly for large $T$, and
$\int_{T_{0}}^{\infty}\frac{1}{\alpha^{0}}\,dT\geq\int_{T_{0}}^{\infty}\frac{N}{\sqrt{f}}\,dT\gtrsim\int_{T_{0}}^{\infty}N\,dT.$
Since $N=n+\mathcal{O}(e^{-T})$, the lapse remains uniformly bounded below by a positive constant, implying
$\int_{T_{0}}^{\infty}N\,dT=\infty.$
It follows that $\lambda(T)\to\infty$ as $T\to\infty$, establishing future completeness for all causal geodesics.
$\square$
Remark 12 (Non-convergence to an Einstein Metric).
We emphasize that in general, the trace-free renormalized Ricci tensor $\mathfrak{T}$ does not decay to zero as $T\to\infty$. This indicates that the limiting metric $\widetilde{g}:=\lim_{T\to\infty}g(T)$ fails to be Einstein, even though it has constant scalar curvature. The underlying obstruction is geometric: the compact manifold MM may not admit any Einstein metric (e.g., hyperbolic metric in the case $n=3$), and hence the evolution governed by the Einstein equations with $\Lambda>0$ does not necessarily drive the geometry toward an Einstein configuration.
To substantiate this, consider the evolution of the squared $L^{2}$-norm of $\mathfrak{T}$,
(201) $S(T):=\|\mathfrak{T}(T)\|_{L^{2}(M)}^{2}.$
Using the evolution equation for $\mathfrak{T}$ and the uniform high-order energy bounds obtained in Section 5.1, along with standard Sobolev inequalities and Grönwall-type arguments, one derives the differential inequality
(202) $\frac{d}{dT}S(T)\geq-C(\mathcal{I}^{0})^{2}e^{-2T},$
where $C>0$ is a constant depending only on the background geometry and the constants in the elliptic-hyperbolic structure of the system. Integrating this inequality from $T_{0}$ to $T$, we obtain
(203) $S(T)\geq S(T_{0})-C(\mathcal{I}^{0})^{2}e^{-2T_{0}}\left(1-e^{-2(T-T_{0})}\right).$
Letting $T\to\infty$, we find
(204) $\liminf_{T\to\infty}\|\mathfrak{T}(T)\|_{L^{2}(M)}^{2}\geq S(T_{0})-C(\mathcal{I}^{0})^{2}e^{-2T_{0}}.$
Hence, for sufficiently large initial energy $\mathcal{I}^{0},T_{0}$, and correspondingly large $S(T_{0})$ (which is permitted by our assumptions), one may ensure that
(205) $\liminf_{T\to\infty}\|\mathfrak{T}(T)\|_{L^{2}(M)}^{2}\gtrsim 1,$
i.e., $\mathfrak{T}$ remains bounded away from zero as $T\to\infty$. This implies that the limiting metric $\widetilde{g}$ fails to satisfy the Einstein condition $\operatorname{Ric}_{\widetilde{g}}=\lambda\widetilde{g}$ for any $\lambda\in\mathbb{R}$, although it satisfies the weaker condition of constant scalar curvature. Notably, the evolution does not asymptote to a hyperbolic geometry even when $M$ admits one. Thus, the asymptotic geometry is generally not Einstein in the presence of a positive cosmological constant.
5.2. Proof of Corollary 1.2
Proof.
Recall that the space of Riemannian metrics on a closed 3-manifold $M$ satisfying the uniform bounds
$|\mathrm{Riem}(g)|\leq C,\quad\operatorname{Vol}(M,g)\geq v>0,\quad\operatorname{diam}(M,g)\leq D,$
for fixed constants $C,v,D>0$, is precompact in the $C^{1,\alpha}$ and $L^{2,p}$ topologies for appropriate $\alpha\in(0,1)$ and $p>3$.
For any $\epsilon>0$, define the $\epsilon$-thick and $\epsilon$-thin parts of $(M,g)$ by
$M^{\epsilon}:=\{x\in M\mid\operatorname{Vol}(B_{x}(1))\geq\epsilon\},\quad M_{\epsilon}:=\{x\in M\mid\operatorname{Vol}(B_{x}(1))<\epsilon\},$
where $B_{x}(1)$ is the geodesic ball of radius 1 centered at $x$ with respect to $g$. The Bishop-Gromov volume comparison theorem [43] implies that for any sequence of metrics $\{g_{i}\}$ satisfying the uniform bounds above, the thin part is empty for sufficiently small $\epsilon$, i.e., $M=M^{\epsilon}$.
By Theorem 1.1, the uniform $L^{2}$-based energy estimates on curvature and the metric components imply uniform control on $|\mathrm{Riem}(g(T))|$, diameter, and volume lower bounds for all $T\geq T_{0}$.
Furthermore, consider the variation of the volume of a fixed geodesic ball $B(1)$ under the CMC evolution:
$\left|\operatorname{Vol}(B(1))_{T}-\operatorname{Vol}(B(1))_{T_{0}}\right|=\left|\int_{T_{0}}^{T}\frac{d}{dT^{\prime}}\operatorname{Vol}(B(1))_{T^{\prime}}\,dT^{\prime}\right|.$
By the first variation formula for volume under the flow, and using the uniform exponential decay estimates for the lapse function $N$ (cf. Section 4.3), we obtain
$\left|\operatorname{Vol}(B(1))_{T}-\operatorname{Vol}(B(1))_{T_{0}}\right|\lesssim\operatorname{Vol}(B(1))_{T_{0}}\int_{T_{0}}^{T}\left\|\frac{N}{n}-1\right\|_{L^{\infty}(B(t^{\prime}))}dt^{\prime}\lesssim e^{-2T_{0}}(1-e^{-2(T-T_{0})}).$
Thus, for sufficiently large initial time $T_{0}$, the local volume $\operatorname{Vol}(B(1))_{T}$ remains uniformly comparable to $\operatorname{Vol}(B(1))_{T_{0}}$, establishing the uniform boundedness claimed in Corollary 1.2. ∎
Theorem (Ringström’s Conjecture in the $\Lambda>0$, $\sigma(M)\leq 0$ Setting).
[…]
$M_{T}\not\subset J^{-}(\gamma).$
Proof.
Let $\gamma(s)=(T(s),x^{i}(s))$ be any future-directed inextendible causal curve. In CMC-transported coordinates, the rescaled spacetime metric takes the form
$\widehat{g}=-N^{2}dT^{2}+g_{ij}(T)dx^{i}dx^{j},$
with lapse $N$ uniformly bounded above by the estimates $\|\frac{N}{n}-1\|_{L^{\infty}(M)}\lesssim e^{-2T},~T\gg 1$. Causality implies
$g_{ij}(T)\frac{dx^{i}}{dT}\frac{dx^{j}}{dT}\leq N^{2}(T)\lesssim|\tau|^{-2}.$
In light of the estimates of theorem 1.1 and the scaling property of 59 in section 3, the physical metric $\widetilde{g}(T)\approx\varphi^{-2}g(T)$ for the re-scaled nonphysical metric $g(T)$. Hence,
$\left\|\frac{dx^{i}}{dT}\right\|_{g(T)}\lesssim\bigg|\frac{\varphi}{\tau}\bigg|.$
Note that for the pure vacuum i.e., $\Lambda=0$, one has $\varphi=\tau$ and therefore the right hand side does not have a decay structure. For $\Lambda>0$, one observes $|\varphi(T)|\lesssim e^{-T}$ and $|\tau|\lesssim\sqrt{e^{-2T}+3\Lambda}=\sqrt{e^{-2T}+3n(n+1)}\gtrsim\sqrt{n(n+1)}$ since we fixed $\Lambda=n(n+1)$ in our framework. The length of the spatial projection $\pi_{M}\circ\gamma$ of $\gamma$ in $(M,g(T))$ satisfies
$\text{Length}_{\bar{g}}[\pi_{M}\circ\gamma]\leq\int_{T}^{\infty}Cn^{-1}(n+1)^{-1}e^{-T^{{}^{\prime}}}dT^{{}^{\prime}}=Cn^{-1}(n+1)^{-1}e^{-T}\lesssim n^{-1}(n+1)^{-1}=\Lambda^{-1}.$
Thus, $\gamma$’s spatial projection $\pi_{M}\circ\gamma$ cannot intersect all of $M_{T}$. Therefore, $M_{T}\not\subset J^{-}(\gamma)$. ∎
6. Funding and/or Conflicts of interests/Competing interests
The authors have no conflicts to disclose.
7. Data Availability
Data sharing is not applicable to this article as no new data were created or analyzed in this study.